### v4

# Open Source Implementations of Numerical Algorithms for

Abstract
The complete elliptic integral of the first kind (CEI-1) plays in a significant role in mathematics, physics and engineering. There is no simple formula for its computation, thus numerical algorithms are essential for coping with practical problems involved. The commercial implementations for the numerical solutions, such as the functions ellipticK and EllipticK provided by MATLAB and Mathematica respectively, are based on ${\mathcal{K}}_{\mathrm{cs}}(m)$ instead of the usual form $K(k)$ such that ${\mathcal{K}}_{\mathrm{cs}}(k^{2})=K(k)$ and $m=k^{2}$. It is necessary to develop open source implementations for the computation of the CEI-1 in order to avoid potential risks of using commercial software and possible limitations due to the unknown factors. In this paper, the infinite series method, arithmetic-geometric mean (AGM) method, Gauss-Chebyshev method and Gauss-Legendre methods are discussed in details with a top-down strategy. The four key algorithms for computing CEI-1 are designed, verified, validated and tested, which can be utilized in R& D and be reused properly. Numerical results show that our open source implementations based on $K(k)$ are equivalent to the commercial implementation based on ${\mathcal{K}}_{\mathrm{cs}}(m)$. The general algorithms for computing orthogonal polynomials developed are significant byproducts in the sense of STEM education and scientific computation.
Keywords: Complete elliptic integral of the first kind (CEI-1); Orthogonal polynomials; Verification-Validation-Testing (VVT); Infinte series; Arithmetic-geometric mean (AGM); Gauss numerical integration
1 Introduction
The Complete Elliptic Integral of the first kind (CEI-1) is an interesting special function in mathematics which also plays a significant role in physics and engineering applications. There are two definitions for the CEI-1: the first one is given by [1, 2, 3]
$K(k)=\int^{\pi/2}_{0}\frac{\operatorname{d}u}{\sqrt{1-k^{2}\sin^{2}u}},\quad k\in[0,1].$ (1)
and the second one is given by
${\mathcal{K}}_{\mathrm{cs}}(m)=\int^{\pi/2}_{0}\frac{\operatorname{d}u}{\sqrt{1-m\sin^{2}u}},\quad m\in[0,1]$ (2)
which is implemented by the commercial software, denoted by the subscript cs of the function ${\mathcal{K}}_{\mathrm{cs}}(\cdot)$, such as the ellipticK function in MATLAB and the EllipticK function in Mathematica. Obviously, the equations (1) and (2) are connected by
[…]
for the arguments $k$ and $m$.
For the special function $K(\cdot)$ defined by integral, the computation should be done with proper numerical algorithm. In practical applications, there are different implementations of numerical algorithms for the CEI-1. Figure 1 illustrates the numerical results obtained by the functions ellipticK in MATLAB and EllipticK in Mathematica. However, the commercial implementations are limited by the licenses. Moreover, the application of MATLAB is prohibited for some affiliations due to some potential conflicts involved in national security.
(a) Output of EllipticK in Mathematica
(b) Output of ellipticK in MATLAB
Figure 1: Value of ${\mathcal{K}}_{\mathrm{cs}}(m)$ obtained by the MATLAB and Mathematica for $m=[0.00,0.01,0.20,0.30,0.40,0.50]$
Actually, we can find some fundamental types of methods to compute the CEI-1 properly:
  • expanding the CEI-1 into an infinite series which converges rapidly;
  • computing the CEI-1 with the arithmetic-geometric mean (AGM) method which converges rapidly and iterates simply;
  • converting the CEI-1 into a finite sum with $12n-$ 1 order algebraic precision by taking $n$-th order orthogonal polynomials via Gauss numerical integration approach.
The Gauss numerical method can be expressed by[4]
$I(f)=\int^{b}_{a}\rho(x)f(x)\operatorname{d}x=\sum^{n}_{i=1}w_{i}f(x_{i})+\mathcal{E}_{n}(f)$ (4)
where $n$ is a positive integer, $\left\{x_{i}:0\leq i\leq n-1\right\}$ is the set of $n$ roots of an orthogonal polynomial, say $G_{n}(x)$, $w_{i}$ are the weights specified by $x_{i}$ and $G_{n}(x)$, and $\mathcal{E}_{n}(f)$ is the error of optimal approximation which will be omitted according to the requirement of precision in practice. However, the integral $K(k)$ is more complicated since it is a function defined by an integral. With the help of the concepts of abstract function defined by an integral, we can use the following form
$\begin{split}K(k)&=\int^{b}_{a}f(x,k)\operatorname{d}\mu=\int^{b}_{a}f(x,k)\rho(x)\operatorname{d}x\\ &=\sum^{n}_{i=1}w_{i}f(x_{i},k)+\mathcal{E}_{n}(f).\end{split}$ (5)
There are two feasible methods in this way:
  • Gauss-Chebyshev method, where $\rho(x)=\frac{1}{\sqrt{1-x^{2}}}$and $b=-a=1$;
  • Gauss-Legendre method, where $\rho(x)=1$, $a=0$ and $b=\pi/2$.
In this paper, we will discuss four methods to calculate the special function $K(k)$ as shown in Figure 2.
Figure 2: Four methods for computing CEI-1
The contents of this paper are organized as following: section 2 deals with the mathematical preliminaries for the infinite series methods, AGM method, Gauss-Chebyshev method and Gauss-Legendre method for computing CEI-1; section 3 copes with the numerical algorithms for the four methods mentioned above; section 4 gives the numerical results for verification and validation of the numerical solvers proposed and finally section 5 gives the conclusions.
2 Mathematical Preliminaries
2.1 Infinite Series Method
The first method for computing the CEI-1 is to take the series expansion of variable $k$. Actually, with the help of the generalized Newton’s binomial theorem, we have [5]
$(1+z)^{\alpha}=\sum^{\infty}_{j=0}\binom{\alpha}{j}z^{j},\quad\left|z\right|\leq 1,\forall\alpha\in\mathbb{R}$ (6)
where
$\binom{\alpha}{0}=1$ (7)
and
$\binom{\alpha}{j}=\frac{\alpha(\alpha-1)\cdots(\alpha-j+1)}{j!}=\prod^{j-1}_{t=0}\frac{\alpha-t}{j-t}$ (8)
is the binomial coefficient. Put $z=-k^{2}\sin^{2}u$ and $\alpha=-\frac{1}{2}$, we can obtain
$\frac{1}{\sqrt{1-k^{2}\sin^{2}u}}=\sum^{\infty}_{j=0}(-1)^{j}\binom{-\frac{1}{2}}{j}k^{2j}\sin^{2j}u.$ (9)
According to the identity
$(-1)^{j}\binom{-r}{j}=\binom{r+j-1}{j},\quad r\in\mathbb{R},j\in\mathbb{N}$ (10)
and integral formulae
$\int^{\pi/2}_{0}\sin^{2j}u\operatorname{d}u=\frac{(2j-1)!!}{(2j)!!}\cdot\frac{\pi}{2},\quad j\in\mathbb{N}$ (11)
where
$\left\{\begin{split}(2j-1)!!&=(2j-1)(2j-3)\cdots 3\cdot 1\\ (2j)!!&=(2j)(2j-2)\cdots 4\cdot 2\\ 0!!&=(-1)!!=1\end{split}\right.$ (12)
are the double factorial. Therefore, we can find the infinite series for the CEI-1, viz.,
$K(k)=\frac{\pi}{2}\cdot\sum^{\infty}_{j=0}\left[\frac{(2j-1)!!}{(2j)!!}\right]^{2}k^{2j}$ (13)
Simple algebraic operations show that
$K(k)=\frac{\pi}{2}\cdot\left[1+\frac{1}{4}k^{2}+\frac{9}{64}k^{4}+\frac{25}{256}k^{6}+\cdots\right]$ (14)
Let
$\quad c_{j}=\left[\frac{(2j-1)!!}{(2j)!!}\right]^{2},\quad j\in\mathbb{N},$ (15)
then it is easy to find that
[…]
according to (3) and (16).
From a computational point of view, we just need to calculate the following convergent infinite series
$S(x)=\sum^{\infty}_{j=0}c_{j}x^{j},\quad x\in[0,1].$ (17)
Therefore, the key for determining CEI-1 lies in designing an algorithm for computing the infinite series (17) for $x=m$ or $x=k^{2}$ and designing a procedure to calculate the coefficient $c_{j}$ specified by (15).
2.2 AGM Method
The second method for calculating the CEI-1 is the AGM method proposed by King [6]. The AGM method was first discovered by Lagrange and then rediscovered and pushed forward greatly by Gauss. The relation of AGM method and CEI-1 was also discussed by Borwein [7] for the purpose of calculating the constant $\pi$. For the 2-dim sequence $\left\{(u_{i},v_{i})\right\}^{\infty}_{j=0}$ such that
$\left\{\begin{split}u_{i+1}&=\frac{u_{i}+v_{i}}{2},\quad i\geq 0\\ v_{i+1}&=\sqrt{u_{i}\cdot v_{i}},\quad i\geq 0\\ u_{0}&=a\\ v_{0}&=b\end{split}\right.$ (18)
it is easy to show that: both the limit of arithmetic mean $u_{i}$ and the limit of geometric mean $v_{i}$ exist, and the limits for $u_{i}$ and $v_{i}$ are equal. Obviously, the common limit just depends on the initial non-negative values $a$ and $b$. This limit specified by the arithmetic-geometric mean (AGM) can be denoted by
$\operatorname{AGM}(a,b)=\lim_{i\to\infty}u_{i}=\lim_{i\to\infty}v_{i}.$ (19)
The updating relation in equation (18) implies that both the symmetric property
$\operatorname{AGM}(a,b)=\operatorname{AGM}(b,a)$ (20)
and the iterative property
$\operatorname{AGM}(a,b)=\operatorname{AGM}\left(\frac{a+b}{2},\sqrt{ab}\right)$ (21)
hold obviously.
There are lots of formulae about the AGM and the CEI-1, we just list the following necessary ones according to our requirements:
$\frac{1}{\operatorname{AGM}(1+x,1-x)}$ $=\sum^{\infty}_{j=0}\left[\frac{(2j-1)!!}{(2j)!!}\right]^{2}x^{2j}$ (22)
$\frac{1}{\operatorname{AGM}(a,b)}$ $=\frac{2}{\pi}\int^{\frac{\pi}{2}}_{0}\frac{\operatorname{d}\theta}{\sqrt{a^{2}\cos^{2}\theta+b^{2}\sin^{2}\theta}}$ (23)
$=\frac{4}{(a+b)\pi}K\left(\left|\frac{a-b}{a+b}\right|\right)$ (24)
For $k\in[0,1],a=1+k$ and $b=1-k$, we can deduce that
$K(k)=\frac{\pi/2}{\operatorname{AGM}(1+k,1-k)}$ (25)
from (22). Consequently, we can obtain
$\operatorname{AGM}(1+k,1-k)=\operatorname{AGM}(1,\sqrt{1-k^{2}})$ (26)
by the iterative property (21). Hence we have
$K(k)=\frac{\pi/2}{\operatorname{AGM}(1+k,1-k)}=\frac{\pi/2}{\operatorname{AGM}(1,\sqrt{1-k^{2}})}$ (27)
This is the famous AGM method proposed by King in 1924. Thus, in the sense of computation, the calculation of CEI-1 is equivalent to computing the function $\operatorname{AGM}(a,b)$ via iterative algorithm and calling it with 2-dim parameter $(a,b)=(1,\sqrt{1-k^{2}})$ or equivalently $(a,b)=(1+k,1-k)$. Moreover, the symmetric property of AGM says that the position of $a$ and $b$ in $\operatorname{AGM}(a,b)$ can be swapped without changing the value.
2.3 Gauss-Chebyshev Method
According to Appendix A.1, for the abstract function
${F}_{\mathrm{C}}(k)=\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x=\int^{1}_{-1}\frac{f(at,k)}{\sqrt{1-t^{2}}}\operatorname{d}t,$ (28)
of the Chebyshev type defined by the integral of ${\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}}$ with parameter $k$, we can computed it by the extended Gauss-Chebyshev formulae of numerical integration. Let ${x}^{\mathrm{gc}}_{i}$ denote the $i$-th root of $T_{n}(x)=\cos(n\arccos x)$, i.e. the $n$-th order Chebyshev polynomial of the first kind, and ${w}^{\mathrm{gc}}_{i}$ denote the $i$-th weight corresponding to ${x}^{\mathrm{gc}}_{i}$, then we have
${x}^{\mathrm{gc}}_{i}=\cos\frac{(2i-1)\pi}{2n},\quad{w}^{\mathrm{gc}}_{i}=\frac{\pi}{n},\quad 1\leq i\leq n$ (29)
and
${F}_{\mathrm{C}}(k)=\sum^{n}_{i=1}{w}^{\mathrm{gc}}_{i}f(a{x}^{\mathrm{gc}}_{i},k)+{\mathcal{E}}^{\mathrm{gc}}_{n}(f,a,k).$ (30)
where
${\mathcal{E}}^{\mathrm{gl}}_{n}(f,a,k)=\frac{2\pi\cdot a^{2n}}{2^{2n}(2n)!}f^{(2n)}(\xi,k),\quad\xi\in(-a,a)$ (31)
is the approximation error.
We now introduce three $n$-dim vectors
$\left\{\begin{split}&{\mathbf{x}}^{\mathrm{gc}}={[{x}^{\mathrm{gc}}_{1},\cdots,{x}^{\mathrm{gc}}_{n}]}^{\textsf{T}},\\ &{\mathbf{w}}^{\mathrm{gc}}={[{w}^{\mathrm{gc}}_{1},\cdots,{w}^{\mathrm{gc}}_{n}]}^{\textsf{T}},\\ &{\mathbf{v}}^{\mathrm{gc}}_{f}(k,a)={[f(a{x}^{\mathrm{gc}}_{1},k),\cdots,f(a{x}^{\mathrm{gc}}_{n},k)]}^{\textsf{T}}\end{split}\right.$ (32)
then CEI-1 can be represented by the inner product of ${\mathbf{w}}^{\mathrm{gc}}$ and ${\mathbf{v}}^{\mathrm{gc}}_{f}(k)$, viz.,
${F}_{\mathrm{C}}(k)=\langle{\mathbf{w}}^{\mathrm{gc}}|{\mathbf{v}}^{\mathrm{gc}}_{f}(k,a)\rangle+{\mathcal{E}}^{\mathrm{gc}}_{n}(f,a,k)$ (33)
in which the inner product is defined by
$\langle\mathbf{w}|\mathbf{v}\rangle=\sum^{n}_{i=1}w_{i}v_{i},\quad\forall\mathbf{w},\mathbf{v}\in\mathbb{R}^{{n}\times{1}}.$ (34)
Our objective here is to give a computational method for CEI-1 by establishing the relation of CEI-1 and ${F}_{\mathrm{C}}(k)$. Let $t=\sin u$, then we have
$\begin{split}K(k)&=\int^{\pi/2}_{0}\frac{\operatorname{d}t}{\sqrt{1-k^{2}\sin^{2}u}}=\int^{1}_{0}\frac{\operatorname{d}t}{\sqrt{(1-t^{2})(1-k^{2}t^{2})}}\\ &=\int^{1}_{-1}\frac{1}{\sqrt{1-t^{2}}}\cdot\frac{1/2}{\sqrt{1-k^{2}t^{2}}}\operatorname{d}t\end{split}$ (35)
Now put $a=1$ and
${f}_{\mathrm{gc}}(t,k)=\frac{1/2}{\sqrt{1-k^{2}t^{2}}},\quad t\in(-1,1),k\in[0,\sqrt{2}/2)$ (36)
then
$\begin{split}K(k)&=\sum^{n}_{i=1}{w}^{\mathrm{gc}}_{i}{f}_{\mathrm{gc}}({x}^{\mathrm{gc}}_{i},k)+{R}^{\mathrm{gc}}_{n}({f}_{\mathrm{gc}})\\ &=\langle{\mathbf{w}}^{\mathrm{gc}}|{\mathbf{v}}^{\mathrm{gc}}_{{f}_{\mathrm{gc}}}({\mathbf{x}}^{\mathrm{gc}},k,1)\rangle+{\mathcal{E}}^{\mathrm{gc}}_{n}({f}_{\mathrm{gc}},1,k)\end{split}$ (37)
Therefore, for any $\varepsilon>0$, the CEI-1 can be computed directly with the inner product of vectors ${\mathbf{w}}^{\mathrm{gc}}$ and ${\mathbf{v}}^{\mathrm{gc}}_{{f}_{\mathrm{gc}}}(k,1)$ if we take a sufficiently large $n$ such that $\left|{\mathcal{E}}^{\mathrm{gc}}_{n}({f}_{\mathrm{gc}},1,k)\right|<\varepsilon$ for approximating the integral with fixed parameter $k$.
2.4 Gauss-Legendre Method
According to the Appendix A.2, the abstract function
${F}_{\mathrm{L}}(k)=\int^{b}_{a}f(x,k)\operatorname{d}x=\frac{b-a}{2}\int^{1}_{-1}f\left(\frac{b+a}{2}+\frac{b-a}{2}t\right)\operatorname{d}t$ (38)
defined by the integral of $f(x,k)$ with parameter $k$ such that $f(\cdot,k)\in C^{2n}[a,b]$ can be computed by the extended Gauss-Legendre formulae of numerical integration. Let ${x}^{\mathrm{gl}}_{i}$ denote the $i$-th root of the $n$-th order Legendre polynomial $P_{n}(x)$ and ${w}^{\mathrm{gl}}_{i}$ denote the $i$-th weight corresponding to ${x}^{\mathrm{gl}}_{i}$ such that
${w}^{\mathrm{gl}}_{i}=\frac{2}{n}\cdot\frac{1}{P_{n-1}({x}^{\mathrm{gl}}_{i})P^{\prime}_{n}({x}^{\mathrm{gl}}_{i})}=\frac{2}{[1-({x}^{\mathrm{gl}}_{i})^{2}]\cdot[P^{\prime}_{n}({x}^{\mathrm{gl}}_{i})]^{2}}$ (39)
where $P^{\prime}_{n}(x)$ is the derivative function of $P_{n}(x)$, then we have
$\begin{split}{F}_{\mathrm{L}}(k)&=\frac{b-a}{2}\sum^{n}_{i=1}{w}^{\mathrm{gl}}_{i}f\left(\frac{b+a+(b-a){x}^{\mathrm{gl}}_{i}}{2},k\right)\\ &\quad+{\mathcal{E}}^{\mathrm{gl}}_{n}(f,a,b,k)\end{split}$ (40)
where
${\mathcal{E}}^{\mathrm{gl}}_{n}(f,a,b,k)=\frac{(n!)^{4}}{[(2n)!]^{3}}\cdot\frac{(b-a)^{2n+1}}{2n+1}f^{(2n)}(\xi,k)$ (41)
is the approximation error in which $\xi\in(a,b)$.
We remark that there are no analytical solutions for the roots of $P_{n}(x)=0$ if $n\geq 4$. Fortunately, it can be solved by the Newton’s method for nonlinear equation. For small $n$, say $n\leq 20$, we can look up some available handbooks of mathematics to get the roots $\left\{{x}^{\mathrm{gl}}_{i}:1\leq i\leq n\right\}$. However, this method is not suggested if the $n$ is unknown or not sure. In this case, a feasible, convenient and effective way is to calculate the roots of $P_{n}(x)$ automatically with proper algorithms implemented via some concrete computer programming language and programs.
Let
$\left\{\begin{split}&{\mathbf{x}}^{\mathrm{gl}}={\left[{x}^{\mathrm{gl}}_{1},\cdots,{x}^{\mathrm{gl}}_{n}\right]}^{\textsf{T}}\\ &{\mathbf{w}}^{\mathrm{gl}}={\left[{w}^{\mathrm{gl}}_{1},\cdots,{w}^{\mathrm{gl}}_{n}\right]}^{\textsf{T}}\\ &{u}^{\mathrm{gl}}_{i}=\frac{b+a+(b-a){x}^{\mathrm{gl}}_{i}}{2},\quad 1\leq i\leq n\\ &{\mathbf{v}}^{\mathrm{gl}}_{f}({\mathbf{x}}^{\mathrm{gl}},k,a,b)=\frac{b-a}{2}{\left[f({u}^{\mathrm{gl}}_{i},k),\cdots,f({u}^{\mathrm{gl}}_{n})\right]}^{\textsf{T}}\end{split}\right.$ (42)
then ${F}_{\mathrm{L}}(k)$ can be represented by the inner product of ${\mathbf{w}}^{\mathrm{gl}}$ and ${\mathbf{v}}^{\mathrm{gl}}_{f}({\mathbf{x}}^{\mathrm{gl}},k,a,b)$, viz.,
${F}_{\mathrm{L}}(k)=\langle{\mathbf{w}}^{\mathrm{gl}}|{\mathbf{v}}^{\mathrm{gl}}_{f}({\mathbf{x}}^{\mathrm{gl}},k,a,b)\rangle+{\mathcal{E}}^{\mathrm{gl}}_{n}(f,a,b,k)$ (43)
Suppose that the root-weight pairs $({x}^{\mathrm{gl}}_{i},{w}^{\mathrm{gl}}_{i})$ for $1\leq i\leq n$ were known, we can establish the relation of ${F}_{\mathrm{L}}(k)$ and CEI-1. For $k\in[0,1)$, let $a=0$, $b=\pi/2$, and
${f}_{\mathrm{gl}}(u,k)=\frac{1}{\sqrt{1-k^{2}\sin^{2}u}}$ (44)
then
$\begin{split}K(k)&=\frac{\pi}{4}\sum^{n}_{i=1}{w}^{\mathrm{gl}}_{i}{f}_{\mathrm{gl}}\left(\frac{\pi}{4}(1+{x}^{\mathrm{gl}}_{i}),k\right)+{\mathcal{E}}^{\mathrm{gl}}_{n}\left({f}_{\mathrm{gl}},0,\frac{\pi}{2},k\right)\\ &=\langle{\mathbf{w}}^{\mathrm{gl}}|{\mathbf{v}}^{\mathrm{gl}}_{{f}_{\mathrm{gl}}}({\mathbf{x}}^{\mathrm{gl}},k,0,\frac{\pi}{2})\rangle+{\mathcal{E}}^{\mathrm{gl}}_{n}\left({f}_{\mathrm{gl}},0,\frac{\pi}{2},k\right)\end{split}$ (45)
For any positive $\varepsilon>0$ and fixed parameter $k\in[0,1)$, we can choose a sufficiently large integer $n$ such that the approximation error is small enough, i.e.,
$\left|{\mathcal{E}}^{\mathrm{gl}}_{n}\left({f}_{\mathrm{gl}},0,\frac{\pi}{2},k\right)\right|<\varepsilon.$
The expression above for CEI-1 means that the essential work left for calculating $K(k)$ is to find the root-weight pairs for the Legendre polynomial $P_{n}(x)$.
Although there are several equivalent definitions for the Legendre polynomial, the best way for computing the Legendre polynomial $P_{n}(x)$ is the following iterative formulae
$\left\{\begin{split}P_{n}(x)&=\frac{2n-1}{n}xP_{n-1}(x)-\frac{n-1}{n}P_{n-2}(x),~{}n\geq 2\\ P_{0}(x)&=1;\\ P_{1}(x)&=x.\end{split}\right.$ (46)
This formulae implies that we can compute $P_{n}(x)$ one by one with an iterative process.
The Newton’s method for solving root of a nonlinear equation $f(x,\alpha)=0$ with a constant parameter $\alpha$ is specified by the iterative process with the form
$\begin{split}x[j+1]&={\mathcal{A}}_{\mathrm{Newton}}(x[j])\\ &=x[j]-\frac{f(x[j],\alpha)}{f^{\prime}(x[j],\alpha)},\quad j=0,1,2,\cdots\end{split}$ (47)
which starts from the initial value (guess of the root) $x[0]=x_{0}$ and ends when
$\left|x[j+1]-x[j]\right|<\varepsilon$ (48)
for sufficiently large $j$ and fixed computational error $\varepsilon$. The updating mapping ${\mathcal{A}}_{\mathrm{Newton}}:x[j]\mapsto x[j+1]$ is a contractive mapping defined by
$\begin{split}{\mathcal{A}}_{\mathrm{Newton}}:\mathbb{R}&\to\mathbb{R}\\ x&\mapsto x-\frac{f(x,\alpha)}{f^{\prime}(x,\alpha)}\end{split}$ (49)
For the Legendre polynomial $P_{n}(x)$, there are $n$ different roots ${x}^{\mathrm{gl}}_{i}$ where $i\in\left\{1,2,\cdots,n\right\}$. The iterative process can be described by
$\left\{\begin{split}{x}^{\mathrm{gl}}_{i}[j+1]&={x}^{\mathrm{gl}}_{i}[j]-\frac{P_{n}({x}^{\mathrm{gl}}_{i}[j])}{P^{\prime}_{n}({x}^{\mathrm{gl}}_{i}[j])},\quad j=0,1,2\cdots\\ {x}^{\mathrm{gl}}_{i}[0]&=\cos\frac{\pi(i+0.75)}{n+0.25},\quad 1\leq i\leq n.\end{split}\right.$ (50)
This iterative process is the basis for designing algorithm for solving the roots of interest. Note that we can set $f(x,\alpha)=P_{n}(x)$ where $\alpha=n$ here.
3 Numerical Algorithms
Different numerical algorithms can be obtained according to the four methods discussed to compute the CEI-1 efficiently, robustly and precisely. In order to broaden the applications of the algorithms, we will generalize some specific constants to corresponding variables.
3.1 Infinite Series Algorithm
The equation (13) and its connection with the general infinite series show that there are some key issues for computing CEI-1:
  • designing an algorithm to compute the general infinite series $S(x)=\sum\limits_{j=0}^{\infty}c_{j}x^{j}$ which consists of four sub-tasks:
    • calculating the coefficient $c_{j}$ for each $j\in\mathbb{N}$,
    • computing the value of the general term $c_{j}x^{j}$ for fixed $x\in\mathbb{R}$,
    • deciding whether the general term $c_{j}x^{j}$ is negligible or not for the sum of the infinite series, and
    • calculating the sum of the not-negligible terms;
  • specifying the parameters involved
    • designing a procedure to compute the coefficient $c_{j}$ in the infinite series of CEI-1, and
    • specifying the variable $x$ according to its connection with the variable $k$, i.e., $x=k^{2}$.
  • calling the procedure for the infinite series and multiplying some constant if necessary.
The procedures for calculating CEI-1 with infinite series method is shown in Figure 3, in which CoefFun is a procedure name (pointer to function or similar entity) used to calling the procedure of computing $c_{j}$.
{forest}
[Kseries [InfinteSeriesSolver [CoefFun[CalcCoefCm]] [IsNegligible] ] ]
Figure 3: Procedures involved in calculating the complete elliptic integral CEI-1 with infinite series method.
There is a practical technique for computing $c_{j}$. The double factorials $(2j)!!$! and $(2j-1)!!$! may be large numbers if $j>10$ and the overflow problem may happen in some computer language if the type of data is integer. However, we can cope with the the value of fraction safely instead of dealing with its denominator and numerator separately. Actually, we have
$\frac{(2j-1)!!}{(2j)!!}=\prod^{j-1}_{t=0}\frac{2(j-t)-1}{2(j-t)}=\prod^{j}_{t=1}\left[1-\frac{0.5}{t}\right],\quad t\geq 1.$ (51)
Thus the fraction $\frac{(2j-1)!!}{(2j)!!}$ can be computed by multiplying $1-0.5/t$ one by one with a simple iteration, which not only reduces the computational complexity but also avoid the overflow problem of the factorial operations. Algorithm 3.1 gives the steps of procedure CalcCoefCm for computing the coefficient $c_{m}=\frac{(2m-1)!!}{(2m)!!}$ by the for-loop construction in computer programming language.
 Algorithm 1 Calculate the coefficient $c_{j}=\left(\frac{(2j-1)!!}{(2j)!!}\right)^{2}$ for the $j$-th term of the infinite series for CEI-1
1:Non-negative integer $j$;
2:The coefficient $c_{j}$ of the infinite series in the function $K(k)=\dfrac{\pi}{2}\sum\limits^{\infty}_{j=0}c_{j}k^{2j}$
3:function CalcCoefCm($j$)
4:     if $j==0$ then
5:         return 1;
6:     end if
7:     $\texttt{prod}\leftarrow 1$;
8:     for $t\in\left\langle 1,2,\cdots,j-1,j\right\rangle$ do
9:         $\texttt{prod}\leftarrow\texttt{prod}\cdot(1-0.5/t)$;
10:     end for
11:     $c_{j}\leftarrow\texttt{prod}\cdot\texttt{prod}$;
12:     return $c_{j}$;
13:end function
For a fixed $x$, the value of the infinite series $S(x)=\sum\limits^{\infty}_{j=0}c_{m}x^{m}$ is defined by
$S(x)=\lim_{i\to\infty}S_{i}(x),\quad S_{i}(x)=\sum^{i-1}_{j=0}c_{j}x^{j}$ (52)
The Cauchy’s criteria of convergence shows that the sequence $\left\{S_{m}(x)\right\}^{\infty}_{j=0}$ converges if and only if
$\forall\varepsilon>0,\quad\left|S_{i}(x)-S_{i-1}(x)\right|=\left|c_{i}x^{i}\right|<\varepsilon.$ (53)
In practice, the relative error relation $\frac{\left|S_{i}(x)-S_{i-1}(x)\right|}{\left|S_{i}(x)\right|}<\varepsilon$ is better than the absolute error relation $\left|S_{i}(x)-S_{i-1}(x)\right|<\varepsilon$. This fact can be used to design the procedure for deciding whether the general term $c_{i}x^{i}$ is negligible or not.
 Algorithm 2 Decide whether the general term $c_{j}x^{j}$ is negligible or not
1:The value of the general term $c_{j}x^{j}$: term, the sum of finite terms: sum, and the value of error (precision) $\varepsilon$ with default value $\varepsilon=10^{-9}$.
2:The logical value TRUE or FALSE
3:function IsNegligible($\texttt{term},\texttt{sum},\varepsilon$)
4:     Set a sufficiently small number: $\delta\leftarrow 10^{-10}$;
5:     if $\left|\texttt{term}\right|/(\left|\texttt{sum}\right|+\delta)<\varepsilon$ then
6:         return TRUE;
7:     else
8:         return FALSE;
9:     end if
10:end function
Algorithm 3.1, named by IsNegligible, is utilized to decide whether the $j$-th term $c_{j}x^{j}$ is important for the sum of infinite series $\sum\limits_{j=0}^{\infty}c_{j}x^{j}$ or not. We remark that in Algorithm 3.1 the small number $\delta$ is used to improve the robustness when $\left|S_{j}(x)\right|$ is near to zero.
The procedure for solving the infinite series $\sum\limits^{\infty}_{j=0}c_{j}x^{j}$ can be designed with a do-while or while-do loop, which is named with InfiniteSeriesSolver and shown in Algorithm 3.1. Obviously, the computation of CEI-1 is treated as a special problem of sovling the infinite series by designing a system with the strategy of top-down.
 Algorithm 3 Calculate the sum of the convergent infinite series $\sum^{\infty}_{j=0}c_{j}x^{j}$ (InfinteSeriesSolver)
1:Pointer to function $\texttt{CoefFun}:\mathbb{N}\to\mathbb{R}$ for computing the coefficient $c_{j}$, variable $x$, and precision $\varepsilon$ with default value $\varepsilon=10^{-9}$.
2:The sum of the infinite series
3:function InfinteSeriesSolver($\texttt{CoefFun},x,\varepsilon$)
4:     Set initial values: $\texttt{sum}\leftarrow 0,\texttt{term}\leftarrow 0,j\leftarrow 0$;
5:     do
6:         $c_{j}\leftarrow\texttt{CoefFun}(j)$;
7:         $\texttt{term}\leftarrow c_{j}\cdot x^{j}$;
[…]
9:         $j\leftarrow j+1$;
10:     while $(~{}!~{}\textsc{IsNeglibible}(\texttt{term},\texttt{sum},\varepsilon))$
11:     return sum;
12:end function
It should be noted that the function $\texttt{CoefFun}:\mathbb{N}\to\mathbb{R}$ is used as an argument for the procedure InfiniteSeries in Algorithm 3.1, which implies that the InfiniteSeries is a high order function and the pointer to function will be involved for the implementation with different computer programming languages. We remark that the pointer to function is a concept in the C programming language and its counterparts in Python and MATLAB are lambda expression and function handle respectively.
Now we can design the procedure for calculating the CEI-1 conveniently. Algorithm 3.1 describes the procedure Kseries for computing CEI-1 with the infinite series method.
 Algorithm 4 Calcualting the function CEI-1 with the infinite series method (Kseries)
1:The variable $k\in[0,1)$, precision $\varepsilon$ with the default value $\varepsilon=10^{-9}$.
2:The value of CEI-1
3:function Kseries($k,\varepsilon$)
[…]
5:     $x\leftarrow k^{2}$; // Attention, please!
[…]
7:     $K\leftarrow\texttt{sum}\cdot\pi/2$;
8:     return $K$;
9:end function
3.2 AGM Algorithm
The key issue of the AGM is to compute the limits $\lim\limits_{i\to\infty}u_{i}$ and $\lim\limits_{i\to\infty}v_{i}$ according to (18). The procedures for calculating CEI-1 with the AGM method is shown in Figure 4, in which Update and Dist are two pointers to functions and will be assigned by AgmUpdate in Algorithm 3.2 and DistEuclid in Algorithm 3.2 respectively.
{forest}
[Kagm [AGM [FixedPointSolver [Update[AgmUpdate]] [Dist[DistEuclid]] ] [$a$] [$b$] [$\varepsilon$] ] ]
Figure 4: Procedures for calculating CEI-1 with AGM method.
Let $\mathbf{x}={[x_{1},x_{2}]}^{\textsf{T}}\in\mathbb{R}^{{2}\times{1}}$, then the AGM method can be converted to a 2-dim fixed-point problem with the iterative formulae
$\mathbf{x}_{i+1}={\mathcal{A}}_{\mathrm{agm}}(\mathbf{x}_{i}),\quad i=0,1,2,\cdots$ (54)
such that $\mathbf{x}_{i}\in\mathbb{R}^{{2}\times{1}}_{+}$ and
${\mathcal{A}}_{\mathrm{agm}}(\mathbf{x})=\begin{bmatrix}(x_{1}+x_{2})/2\\ \sqrt{x_{1}x_{2}}\end{bmatrix},\quad\mathbf{x}_{0}=\begin{bmatrix}a\\ b\end{bmatrix}$ (55)
for non-negative $a$ and $b$. The Banach’s fixed-point theorem in functional analysis shows that (54) is convergent when ${\mathcal{A}}_{\mathrm{agm}}$ is contractive ($\left\lVert{\mathcal{A}}_{\mathrm{agm}}\right\rVert<1$). Without doubt, (55) satisfies the convergent condition for any non-negative components $a$ and $b$ of the initial vector $\mathbf{x}_{0}$.
The updating process in AGM method gives a typical nonlinear mapping which is described by the procedure AgmUpdate in Algorithm 3.2.
 Algorithm 5 AGM updating (AgmUpdate)
1:2-dim vector $\mathbf{x}={[x_{1},x_{2}]}^{\textsf{T}}$
2:2-dim vector $\mathbf{y}=\mathcal{A}(\mathbf{x})$
3:function AgmUpdate($\mathbf{x}$)
4:     $y_{1}\leftarrow(x_{1}+x_{2})/2$;
5:     $y_{2}\leftarrow\sqrt{x_{1}x_{2}}$;
6:     $\mathbf{y}\leftarrow{[y_{1},y_{2}]}^{\textsf{T}}$;
7:     return $\mathbf{y}$;
8:end function
It is wise to design a general algorithm for solving the fixed-point in a general way since the fixed-point may be frequently encountered in different STEM problems. The computation of fixed-point is based on the concept of norm. In consequence, it is necessary to design the norm for the vector $\mathbf{x}$. Usually, the Euclidean norm $\left\lVert\mathbf{x}\right\rVert=\sqrt{\sum\limits^{n}_{j=1}x^{2}_{j}}$ for $\mathbf{x}={[x_{1},\cdots,x_{n}]}^{\textsf{T}}\in\mathbb{R}^{{n}\times{1}}$ will be a good choice and the induced 2-dim Euclidean distance will be
$\operatorname{\mathrm{dist}}(\mathbf{x},\mathbf{y})={\left\lVert\mathbf{x}-\mathbf{y}\right\rVert}_{2}=\sqrt{\sum^{n}_{j=1}(x_{j}-y_{j})^{2}}$ (56)
for $\mathbf{x}={[x_{1},\cdots,x_{n}]}^{\textsf{T}}$ and $\mathbf{y}={[y_{1},\cdots,y_{n}]}^{\textsf{T}}$. The procedure DistEuclid described in Algorithm 3.2 is used to compute the $n$-dim Euclidean distance.
 Algorithm 6 Calculate the Euclidean distance of two $n$-dim vectors (DistEuclid).
1:Two $n$-dim vectors $\mathbf{x},\mathbf{y}\in\mathbb{R}^{{n}\times{1}}$.
2:The distance between $\mathbf{x}$ and $\mathbf{y}$.
3:function DistEuclid($\mathbf{x},\mathbf{y},n$)
4:     $\texttt{sum}\leftarrow 0$ ;
5:     for $i\in\left\{1,\cdots,n\right\}$ do
6:         $\texttt{sum}\leftarrow\texttt{sum}+(x_{i}-y_{i})^{2}$;
7:     end for
8:     $\texttt{dist}\leftarrow\sqrt{\texttt{sum}}$;
9:     return dist;
10:end function
The iterative process of AGM implies a special kind of iteration of fixed-point, which can be computed in a generic way with the help of while-do or do-while loop. The procedure FixedPointSolver, as shown in Algorithm 3.2, computes the $n$-dim fixed-point iteratively. Note that the pointers to functions, viz. Update and Dist should be assigned with concrete procedures according to the data type of $\mathbf{x}$.
 Algorithm 7 Compute the abstract fixed-point iteratively (AbstractFixedPointSolver)
1:Dimension $n$, initial data $\mathbf{x}_{0}$, precision $\varepsilon$, updating function Update for the contractive mapping $\mathcal{A}:\mathscr{V}\to\mathscr{V}$, and distance function $\texttt{Dist}:\mathscr{V}\times\mathscr{V}\to\mathbb{R}^{+}$.
2:The fixed-point $\mathbf{x}$ such that $\mathbf{x}=\mathcal{A}(\mathbf{x})$.
3:function AbstractFixedPointSolver($n$, Update, Dist, $\mathbf{x}_{0}$, $\varepsilon$)
4:     $\texttt{guess}\leftarrow\mathbf{x}_{0}$;
5:     $\texttt{improve}\leftarrow\texttt{Update}(\mathbf{x}_{0})$;
6:     while $(\texttt{Dist}(\texttt{guess},\texttt{improve},n)\geq\varepsilon)$ do
[…]
9:     end while
10:     return improve;
11:end function
If the general Banach space $\mathscr{V}$ is just the $n$-dim Euclidean space $\mathbb{R}^{{n}\times{1}}$ and the distance is just the Euclidean distance, then the computation of distance will be easy. Consequently, we can design a procedure FixedPointSolver for computing the fixed-point in $\mathbb{R}^{{n}\times{1}}$. For more details, please see Algorithm 3.2.
For the AGM problem concerned, we have the dimension $n=2$ and the 2-dim vectors can be represented by a 1-dim array of size $2$ or struct with two components when coding with some concrete computer programming language. Therefore, the AGM method can be implemented by taking the procedure FixedPointSolver in Algorithm 3.2 and assigning the pointers to function Update with the procedures AgmUpdate in Algorithm 3.2. Once the iteration ends, we can return the arithmetic mean or the geometric mean in the 2-dim vector according to (19). For more details, please see the procedure AGM in Algorithm 3.2.
 Algorithm 8 Compute the $n$-dim fixed-point $\mathbf{x}\in\mathbb{R}^{{n}\times{1}}$ such that $\mathbf{x}=\mathcal{A}(\mathbf{x})$ with iterative method (FixedPointSolver).
1:Dimension $n$, initial data $\mathbf{x}_{0}$, precision $\varepsilon$, updating function Update for the contractive mapping $\mathcal{A}:\mathbb{R}^{{n}\times{1}}\to\mathbb{R}^{{n}\times{1}}$.
2:The $n$-dim fixed-point $\mathbf{x}\in\mathbb{R}^{{n}\times{1}}$ such that $\mathbf{x}=\mathcal{A}(\mathbf{x})$.
3:function FixedPointSolver($n$, Update, $\mathbf{x}_{0}$, $\varepsilon$)
4:     $\texttt{guess}\leftarrow\mathbf{x}_{0}$;
5:     $\texttt{improve}\leftarrow\texttt{Update}(\mathbf{x}_{0})$;
6:     while $(\textsc{DistEuclid}(\texttt{guess},\texttt{improve},n)\geq\varepsilon)$ do
[…]
9:     end while
10:     return improve;
11:end function
 Algorithm 9 Calculate the value of the arithmetic-geometric mean of non-negative values $a$ and $b$ (AGM).
1:Non-negative values $a$ and $b$ and precision $\varepsilon$.
2:The value of the AGM for $a$ and $b$.
3:function AGM($a,b,\varepsilon$)
4:     $\mathbf{x}_{0}\leftarrow{[a,b]}^{\textsf{T}}$;
[…]
7:     $\texttt{dim}\leftarrow 2$;
8:     $\mathbf{x}\leftarrow\textsc{FixedPointSolver}(\texttt{dim},\texttt{Update},\texttt{Dist},\mathbf{x}_{0},\varepsilon)$;
9:     return $x_{1}$; // return one component of $\mathbf{x}$;
10:end function
The procedure for computing CEI-1 with the AGM method, named by Kagm, is illustrated in Algorithm  3.2 in which the procedure AGM is called directly by setting proper parameters $a\leftarrow 1$ and $b\leftarrow\sqrt{1-k^{2}}$.
 Algorithm 10 Calculate the CEI-1 with the AGM method (Kagm)
1:The variable $k\in[0,1)$, precision $\varepsilon$ with default value $\varepsilon=10^{-9}$.
2:The value of CEI-1
3:function Kagm($k,\varepsilon$)
4:     $a\leftarrow 1$;
5:     $b\leftarrow\sqrt{1-k^{2}}$;
6:     $K\leftarrow(\pi/2)/\textsc{AGM}(a,b,\varepsilon)$;
7:     return $K$;
8:end function
3.3 Gauss-Chebyshev Algorithm
The algorithm of Gauss-Chebyshev for computing CEI-1 is relatively simple due to the equations (29), (36) and (37). Figure 5 demonstrates the procedures involved in calculating CEI-1 with the Gauss-Chebyshev method.
{forest}
[Kintgc [IntegralSolverGC [f[KerGC]] [InnerProduct [${\mathbf{w}}^{\mathrm{gc}}$ [CalcWeightsGC] ] [${\mathbf{v}}^{\mathrm{gc}}_{f}$ [CalcVectorVfGC [${\mathbf{x}}^{\mathrm{gc}}$ [CalcRootsChebyshev] ] [f] [$n$] [$k$] [$a$] ] ] [$n$] ] ] ]
Figure 5: Procedures involved in calculating complete elliptic integral CEI-1 with Gauss-Chebyshev method.
The computation of CEI-1 is recognized as a special problem of solving the general numerical integration ${F}_{\mathrm{C}}(k)$ according to (33), which corresponds to the procedure IntegralSolverGC in Algorithm 3.3, by setting the pointer to function f with KerGC in Algorithm 3.3.
We now give some interpretation for designing the algorithm:
  • The procedure KerGC in Algorithm 3.3 is defined to compute the function ${f}_{\mathrm{gc}}(t,k)=\frac{1/2}{\sqrt{1-k^{2}t^{2}}}$ and use a pointer to function KerFun to represent the ${f}_{\mathrm{gc}}$ in the design and implementation of the algorithm.
  • The root vector ${\mathbf{x}}^{\mathrm{gc}}$ is set by the procedure CalcRootschebyshev in Algorithm 3.3.
  • The value of integral is computed by the procedure InnerProduct in Algorithm 3.3.
  • The arguments of Algorithm 3.3, namely the vector ${\mathbf{w}}^{\mathrm{gc}}$ and vector ${\mathbf{v}}^{\mathrm{gc}}_{f}({\mathbf{x}}^{\mathrm{gc}},k,a)$, are calculated by the procedures CalcWeightsGC in Algorithm 3.3 and CalcVectorVfGC in Algorithm 3.3 respectively.
 Algorithm 11 Calcualte the integral (abstract function)
${F}_{\mathrm{C}}(y)=\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x=\langle{\mathbf{w}}^{\mathrm{gc}}|{\mathbf{v}}^{\mathrm{gc}}_{f}({\mathbf{x}}^{\mathrm{gc}},k,a)\rangle+\mathcal{E}_{n}(f)$
with the parameter $k$ via the $n$-nodes Gauss-Chebyshev numerical integration method (IntegralSolverGC).
1:Variable $k$, pointer to function $\texttt{f}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, positive integral $n$, $n$-dim root vector ${\mathbf{x}}^{\mathrm{gc}}$, $n$-dim weight vector ${\mathbf{w}}^{\mathrm{gc}}$, positive $a$.
2:The value of ${F}_{\mathrm{C}}(y)$.
3:function IntegralSolverGC($k,\texttt{f},n,{\mathbf{x}}^{\mathrm{gc}},{\mathbf{w}}^{\mathrm{gc}},a$)
4:     ${\mathbf{v}}^{\mathrm{gc}}\leftarrow\textsc{CalcVectorVfGC}(\texttt{f},n,{\mathbf{x}}^{\mathrm{gc}},k,a)$;
5:     $\texttt{j}\leftarrow\textsc{InnerProduct}({\mathbf{w}}^{\mathrm{gc}},{\mathbf{v}}^{\mathrm{gc}},n)$;
6:     return j;
7:end function
 Algorithm 12 Compute the function ${f}_{\mathrm{gc}}(t,k)$ in the Gauss-Chebyshev method for calculating the value CEI-1 (KerGC)
1:Variables $t\in[-1,1],k\in[0,1)$
2:The value of ${f}_{\mathrm{gc}}(t,k)$
3:function KerGC($t,k$)
4:     $u\leftarrow kt$;
5:     $v\leftarrow 0.5/\sqrt{1-u^{2}}$;
6:     return $v$;
7:end function
 Algorithm 13 Calculate the root vector for the $n$ roots of the Chebyshev polynomial $T_{n}(x)=\cos(n\arccos x)$ (CalcRootsChebyshev)
1:Positive integer $n$.
2:The $n$ roots of $T_{n}(x)$.
3:function CalcRootsChebyshev($n$)
4:     for $i\in\left\{1,\cdots,n\right\}$ do
5:         ${x}^{\mathrm{gc}}_{i}\leftarrow\cos\cfrac{(2i-1)\pi}{2n}$;
6:     end for
7:     return ${\mathbf{x}}^{\mathrm{gc}}$;
8:end function
 Algorithm 14 Compute the inner product of two $n$-dim vectors. (InnerProduct)
1:Two vectors $\mathbf{w}$ and $\mathbf{v}$, dimension $n$ of the vectors.
2:Inner product $\langle\mathbf{w}|\mathbf{v}\rangle=\sum\limits^{n}_{i=1}w_{i}v_{i}$.
3:function InnerProduct($\mathbf{w},\mathbf{v},n$)
4:     $\texttt{sum}\leftarrow 0$;
5:     for $i\in\left\{1,\cdots,n\right\}$ do
6:         $\texttt{sum}\leftarrow\texttt{sum}+w_{i}\cdot v_{i}$;
7:     end for
8:     return sum;
9:end function
 Algorithm 15 Calculate the weight vector ${\mathbf{w}}^{\mathrm{gc}}$ for the abstract function (parametric integral with parameter $k$) ${F}_{\mathrm{C}}(k)=\displaystyle{\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x}=\langle{\mathbf{w}}^{\mathrm{gc}}|{\mathbf{v}}^{\mathrm{gc}}_{f}({\mathbf{x}}^{\mathrm{gc}},k,a)\rangle+\mathcal{E}_{n}(f)$ with the $n$-nodes Gauss-Chebyshev numerical integration method (CalcWeightsGC)
1:Positive integer $n$
2:The weight vector ${\mathbf{w}}^{\mathrm{gc}}$
3:function CalcWeightsGC($n$)
4:     for $i\in\left\{1,\cdots,n\right\}$ do
5:         ${w}^{\mathrm{gc}}_{i}\leftarrow\pi/n$;
6:     end for
7:     return ${\mathbf{w}}^{\mathrm{gc}}$;
8:end function
 Algorithm 16 Calculate the vector ${\mathbf{v}}^{\mathrm{gc}}_{f}({\mathbf{x}}^{\mathrm{gc}},k,a)$ of function $f(x,k)$ at the x-nodes with the coordinates ${u}^{\mathrm{gc}}_{i}=a{x}^{\mathrm{gc}}_{i}$ (CalcVectorVfGC).
1:Pointer to function $\texttt{f}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, positive integer $n$, $n$-dim root vector ${\mathbf{x}}^{\mathrm{gc}}$, positive $a$, and variable $k$.
2:The vector ${\mathbf{v}}^{\mathrm{gc}}_{f}(k,a)$.
3:function CalcVectorVfGC(f, $n$, ${\mathbf{x}}^{\mathrm{gc}}$, $k$, $a$)
4:     for $i\in\left\{1,\cdots,n\right\}$ do
5:         ${v}^{\mathrm{gc}}_{i}\leftarrow\texttt{f}(a\cdot{x}^{\mathrm{gc}}_{i},k)$ ;
6:     end for
7:     return ${\mathbf{v}}^{\mathrm{gc}}$;
8:end function
 Algorithm 17 Calculate the CEI-1 with the Gauss-Chebyshev method (Kintgc)
1:Variable $k\in[0,1)$, positive integer $n$ with default value $n=50$
2:The value of CEI-1
3:function Kintgc($k,n$)
4:     ${\mathbf{x}}^{\mathrm{gc}}\leftarrow\textsc{CalcRootsChebyshev}(n)$;
5:     ${\mathbf{w}}^{\mathrm{gc}}\leftarrow\textsc{CalcWeightsGC}({\mathbf{x}}^{\mathrm{gc}},n)$;
[…]
7:     $a\leftarrow 1.0$;
8:     $K\leftarrow\textsc{IntegralSolverGC}(k,\texttt{f},n,{\mathbf{x}}^{\mathrm{gc}},{\mathbf{w}}^{\mathrm{gc}},a)$;
9:     return $K$;
10:end function
3.4 Gauss-Legendre Algorithm
The key idea of the Gauss-Legendre method is the same as that of the Gauss-Chebyshev. However, it is more complex due to the following reasons:
  • it is necessary for us to compute the Legendre polynomials $P_{n}(x)$ and their derivatives $P^{\prime}_{n}(x)$ automatically;
  • the root vector ${\mathbf{x}}^{\mathrm{gl}}={\left[{x}^{\mathrm{gl}}_{1},\cdots,{x}^{\mathrm{gl}}_{n}\right]}^{\textsf{T}}$ of $P_{n}(x)$ must be solved iteratively with the Newton’s method;
  • the weight vector ${\mathbf{w}}^{\mathrm{gl}}={\left[{w}^{\mathrm{gl}}_{1},\cdots,{w}^{\mathrm{gl}}_{n}\right]}^{\textsf{T}}$ must be solved with the help of ${\mathbf{x}}^{\mathrm{gl}}$, $P_{n}(x)$ and $P^{\prime}_{n}(x)$;
  • the Newton’s method for solving roots of nonlinear equation should be designed.
The strategy for design algorithm for CEI-1 with Gauss-Legendre method is still top-down. In other words, we first design an algorithm to the general problem of computing the integral ${\int^{b}_{a}f(x,k)\operatorname{d}x}=\langle{\mathbf{w}}^{\mathrm{gl}}|{\mathbf{v}}^{\mathrm{gl}}_{f}({\mathbf{x}}^{\mathrm{gc}},k,a,b)\rangle+\mathcal{E}_{n}(f)$ with inner product, then we calculate the vectors ${\mathbf{x}}^{\mathrm{gl}}$ and ${\mathbf{w}}^{\mathrm{gl}}$ with orthogonal polynomials and Newton’s method for solving roots.
There is also an alternative way to get the root vector ${\mathbf{x}}^{\mathrm{gl}}$ and ${\mathbf{w}}^{\mathrm{gl}}$ by reading some handbooks about special functions and/or numerical analysis. However, the tables obtained are just for some fixed significant digits and fixed order $n$ such that $n\leq 20$ for $P_{n}(x)$. It is necessary for us to point out that a smart way born in 1900s is not always a wise way in 2020s. Moreover, the tables given in references can be used to check whether the output of the implementation of our algorithms is correct or not. It is a good idea that we respect the tradition, stand on the tradition, but also try to go beyond the tradition.
The complete design for the algorithm of Gauss-Legendre for CEI-1 can be divided into four sub-tasks. We now discuss them one-by-one with a strategy of top-down. Figure 6 demonstrates the structure of procedures in the Gauss-Legendre method.
{forest}
[Kintgl [IntegralSolverGL [f[KerGL]] [InnerProduct [${\mathbf{w}}^{\mathrm{gl}}$[CalcWeightsGL]] [${\mathbf{v}}^{\mathrm{gl}}_{f}$ [CalcVectorVfGL [${\mathbf{x}}^{\mathrm{gl}}$ [CalcRootsLegendre] ] [f] [$n$] [$k$] [$a$] [$b$] ] ] [$n$] ] ] ]
Figure 6: Procedures involved in calculating the complete elliptic integral CEI-1 with Gauss-Legendre method.
3.5 Compute CEI-1 with the Gauss-Legendre based on the roots, weights and kernel function
Firstly, recognize the computation of CEI-1 as a special problem of computing the integral ${F}_{\mathrm{L}}(k)$ with numerical integration via (44) and (45). The procedure IntegralSolverGL in Algorithm 3.5 is used to compute ${F}_{\mathrm{L}}(k)$ based on the procedure InnerProduct in Algorithm 3.3 and the procedure KerGL in Algorithm 3.5 to compute the function ${f}_{\mathrm{gl}}(t,k)=1/\sqrt{1-k^{2}\sin^{2}t}$.
 Algorithm 18 Calcualte the parametric integral
${F}_{\mathrm{L}}(y)=\int^{b}_{a}f(x,k)\operatorname{d}x=\langle{\mathbf{w}}^{\mathrm{gl}}|{\mathbf{v}}^{\mathrm{gl}}_{f}({\mathbf{x}}^{\mathrm{gl}},k,a,b)\rangle+\mathcal{E}_{n}(f)$
via the $n$-nodes Gauss-Legendre numerical integration method (IntegralSolverGL).
1:Variable $k$, pointer to function $\texttt{f}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, positive integral $n$, $n$-dim root vector ${\mathbf{x}}^{\mathrm{gl}}$, $n$-dim weight vector ${\mathbf{w}}^{\mathrm{gl}}$, variable $a$ and variable $b$.
2:The value of ${F}_{\mathrm{L}}(y)$.
3:function IntegralSolverGL($k$, f, $n$, ${\mathbf{x}}^{\mathrm{gl}}$, ${\mathbf{w}}^{\mathrm{gl}}$, $a$, $b$)
4:     ${\mathbf{v}}^{\mathrm{gl}}\leftarrow\textsc{CalcVectorVfGL}(\texttt{f},n,{\mathbf{x}}^{\mathrm{gl}},k,a,b)$;
5:     $\texttt{j}\leftarrow\textsc{InnerProduct}({\mathbf{w}}^{\mathrm{gl}},{\mathbf{v}}^{\mathrm{gl}},n)$;
6:     return j;
7:end function
 Algorithm 19 Compute the function ${f}_{\mathrm{gl}}(t,k)$ in the Gauss-Legendre method for calculating the value CEI-1 (KerGL)
1:Variables $t\in[-1,1],k\in[0,1)$
2:The value of ${f}_{\mathrm{gl}}(t,k)$
3:function KerGL($t,k$)
4:     $u\leftarrow k\cdot\sin t$;
5:     $v\leftarrow 1/\sqrt{1-u^{2}}$;
6:     return $v$;
7:end function
 Algorithm 20 Calculate the vector ${\mathbf{v}}^{\mathrm{gl}}_{f}({\mathbf{x}}^{\mathrm{gl}},k,a,b)$ of function $f(x,k)$ at the x-nodes with the coordinates ${u}^{\mathrm{gl}}_{i}=\left[b+a+(b-a){x}^{\mathrm{gc}}_{i}\right]/2$ (CalcVectorVfGL).
1:Pointer to function $\texttt{f}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, order $n$, $n$-dim root vector ${\mathbf{x}}^{\mathrm{gl}}$, variable $a$, variable $b$, and variable $k$.
2:Vector ${\mathbf{v}}^{\mathrm{gc}}_{f}({\mathbf{x}}^{\mathrm{gl}},k,a,b)$.
3:function CalcVectorVfGL(f, $n$, ${\mathbf{x}}^{\mathrm{gc}}$, $k$, $a$, $b$)
4:     for $i\in\left\{1,\cdots,n\right\}$ do  //${v}^{\mathrm{gl}}_{f,i}({\mathbf{x}}^{\mathrm{gl}},k,a,b)$;
5:         ${u}^{\mathrm{gl}}\leftarrow(b+a+(b-a){x}^{\mathrm{gc}}_{i})/2$;
6:         ${v}^{\mathrm{gl}}_{i}\leftarrow\texttt{f}({u}^{\mathrm{gl}},k)\cdot\frac{b-a}{2}$;
7:     end for
8:     return ${\mathbf{v}}^{\mathrm{gl}}$;
9:end function
 Algorithm 21 Calculate the CEI-1 with the Gauss-Legendre method (Kintgl)
1:The variable $k\in[0,1)$, positive integer $n$ with default value $n=50$
2:The value of CEI-1
3:function Kintgl($k,n$)
4:     ${\mathbf{x}}^{\mathrm{gl}}\leftarrow\textsc{CalcRootsLegendre}(n)$;
5:     ${\mathbf{w}}^{\mathrm{gl}}\leftarrow\textsc{CalcWeightsGL}({\mathbf{x}}^{\mathrm{gl}},n)$;
[…]
7:     $a\leftarrow 0$;
8:     $b\leftarrow\pi/2$;
9:     $K\leftarrow\textsc{IntegralSolverGL}(k,f,n,{\mathbf{x}}^{\mathrm{gl}},{\mathbf{w}}^{\mathrm{gl}},a,b)$;
10:     return $K$;
11:end function
3.6 Generating the Weights Specified by the Legendre polynomial and its Derivative
The weight vector ${\mathbf{w}}^{\mathrm{gl}}$ for the Gauss-Legendre numerical integration can be generated with the help of the root vector ${\mathbf{x}}^{\mathrm{gl}}$. According to (39), we design a procedure CalcWeightsGL in Algorithm 3.6 to calculate ${\mathbf{w}}^{\mathrm{gl}}$ by calling the procedure DerivPolynomLegendre in Algorithm 3.8. The details for the procedure CalcWeightsGL are given in Algorithm  3.6, which is a direct translation of (39) from mathematics to pseudo-code. The structure of procedures for calculating the weights is shown in Figure 7.
{forest}
[CalcWeightsGL [${\mathbf{x}}^{\mathrm{gc}}$[CalcRootsLegendre]] [$P^{\prime}_{n}(x)$[DerivPolynomLegendre]] ]
Figure 7: Procedures involved in computing the $n$ weights for the numerical integration with Gauss-Legendre method.
 Algorithm 22 Calculate the weight vector ${\mathbf{w}}^{\mathrm{gl}}$ for the abstract function (parametric integral with parameter $k$) ${F}_{\mathrm{L}}(k)=\displaystyle{\int^{b}_{a}f(x,k)\operatorname{d}x}=\langle{\mathbf{w}}^{\mathrm{gl}}|{\mathbf{v}}^{\mathrm{gl}}_{f}({\mathbf{x}}^{\mathrm{gl}},k,a,b)\rangle+\mathcal{E}_{n}(f)$ with the $n$-nodes Gauss-Legendre numerical integration method (CalcWeightsGL).
1:Positive integer $n$ for the number of weights, root vector ${\mathbf{x}}^{\mathrm{gl}}$ of $P_{n}(x)$.
2:Weight vector ${\mathbf{w}}^{\mathrm{gl}}$ for the Gauss-Legendre numerical integration.
3:function CalcWeightsGL(${\mathbf{x}}^{\mathrm{gl}},n$)
4:     for $i\in\left\{1,\cdots,n\right\}$ do
5:         $u\leftarrow\textsc{DerivPolynomLegendre}({x}^{\mathrm{gl}}_{i},n)$;
6:         ${w}^{\mathrm{gl}}_{i}\leftarrow\cfrac{2}{[1-({x}^{\mathrm{gl}}_{i})^{2}]u^{2}}$;
7:     end for
8:     return ${\mathbf{w}}^{\mathrm{gl}}$;
9:end function
3.7 Solving the $n$ Roots of the Legendre Polynomial $P_{n}(x)$
Figure 8: Solving the root vector ${\mathbf{x}}^{\mathrm{gl}}$ for the Legendre polynomial $P_{n}(x)$ such that $P_{n}({x}^{\mathrm{gl}}_{j})=0$.
Figure 8 demonstrates the system of solving the root vector of $P_{n}(x)$. Usually in the books about special mathematical functions [8, 9, 10], the different kinds of orthogonal polynomials, such as the Legendre polynomials $P_{n}(x)$, Chebyshev polynomials $T_{n}(x)$ of the first or second kind, Laguerre polynomials $L_{n}(x)$ and Hermite polynomials $H_{n}(x)$, are treated separately, which hides the general iterative formulae (58). Once the $i$-th root ${x}^{\mathrm{gl}}_{i}$ is obtained, the corresponding weight ${w}^{\mathrm{gl}}_{i}$ can be determined by (39).
The steps for solving the $n$ roots of $P_{n}(x)$ lie in three issues:
  • designing the update procedure NewtonUpdate in Algorithm 3.7 by using $P_{n}(x)$ with the iteration formulae in Newton’s method;
  • setting proper initial values ${x}^{\mathrm{gl}}_{i}[0]$ for the $n$ roots of $P_{n}(x)$ according to (50) and call the procedure CalcRootsLegendre in Algorithm 3.7;
  • computing one root of $P_{n}(x)$ by the Newton’s method encapsulated by the procedure RootSolverOrthoPolynom in Algorithm 3.7.
The structure of procedures for calculating the roots is illustrated in Figure 9.
{forest}
[CalcRootsLegendre [RootSolverOrthoPolynom [IsGoodEnough[DistEuclid]] [NewtonUpdate [Df[DerivPolynomLegendre]] [f[PolynomLegendre]] ] ] ]
Figure 9: Procedures involved in calculating the $n$ roots of $P_{n}(x)$.
 Algorithm 23 Calculte the root of orthogonal polynomial $Y_{n}(x)$ with an initial value $x_{0}$ via Newton’s method (RootSolverOrthoPolynom).
1:Pointer to function $\texttt{f}:\mathbb{R}\times\mathbb{N}\to\mathbb{R}$ for $Y_{n}(x)$, pointer to function $\texttt{Df}:\mathbb{R}\times\mathbb{N}\to\mathbb{R}$ for the derivative function $Y^{\prime}_{n}(x)$, initial value $x_{0}$, order $n$, and the precision $\varepsilon$.
2:The root $x$ of $Y_{n}(x)=0$ iterated from $x_{0}$.
3:function RootSolverOrthoPolynom(f, Df, $x_{0}$, $n$, $\varepsilon$)
4:     $\texttt{guess}\leftarrow 0$ ;
5:     $\texttt{improve}\leftarrow x_{0}$;
6:     do
7:         guess $\leftarrow$ improve;
8:         improve $\leftarrow$ NewtonUpdate(f, Df, guess, $n$);
9:     while $(\textsc{DistEuclid}(\texttt{guess},\texttt{improve},1)\geq\varepsilon)$
10:     return improve;
11:end function
 Algorithm 24 Updating function ${\mathcal{A}}_{\mathrm{Newton}}$ for solving a root of the orthogonal polynomial $Y_{n}(x)$ with the Newton method (NewtonUpdate).
1:Two pointers to function: f and Df for $Y_{n}(x)$ and its derivative $Y^{\prime}_{n}(x)$ respectively, variable $x$ and order $n\in\mathbb{N}$.
2:Updated value of the approximation for the root of interest
3:function NewtonUpdate(f, Df, $x$, $n$)
4:     $\texttt{slope}\leftarrow\texttt{Df}(x,n)$; // Compute $Y^{\prime}_{n}(x)$
5:     CheckSlope(slope); // Check $Y^{\prime}_{n}(x)$;
6:     ${x}_{\mathrm{new}}\leftarrow x-\cfrac{\texttt{f}(x,n)}{\texttt{slope}}$; // $x_{i+1}=x_{i}-\cfrac{Y_{n}(x_{i})}{Y^{\prime}_{n}(x_{i})}$
7:     return ${x}_{\mathrm{new}}$;
8:end function
 Algorithm 25 Check whether the slope/derivative at point $x$ is zero or not (CheckSlope).
1:The slope (value of derivative function) slope
2:Void or exception information if necessary
3:function CheckSlope(slope)
4:     $\delta\leftarrow 10^{-10}$;
5:     if ($\left|\texttt{slope}\right|<\delta$then
6:         Print Exception Information!
7:         Exit;
8:     end if
9:end function
 Algorithm 26 Calculate the $n$-dim root vector ${\mathbf{x}}^{\mathrm{gl}}={[{x}^{\mathrm{gl}}_{1},\cdots,{x}^{\mathrm{gl}}_{n}]}^{\textsf{T}}$ of the $n$–th order Legendre polynomial $P_{n}(x)$ such that $P_{n}({x}^{\mathrm{gl}}_{i})=0$ for $1\leq i\leq n$ (CalcRootsLegendre).
1:Positive integer $n$
2:The root vector ${\mathbf{x}}^{\mathrm{gl}}$ of $P_{n}(x)$
3:function CalcRootsLegendre($n$)
4:     Allocate memory for ${\mathbf{x}}^{\mathrm{gl}}={[{x}^{\mathrm{gl}}_{1},\cdots,{x}^{\mathrm{gl}}_{n}]}^{\textsf{T}}$;
5:     $\varepsilon\leftarrow 10^{-9}$; //set the precision for root
[…]
8:     for $i\in\left\{1,2,\cdots,n\right\}$ do
9:         $x_{0}\leftarrow\cos\cfrac{\pi(i+0.75)}{n+0.25}$; //initial value;
10:         ${x}^{\mathrm{gl}}_{i}\leftarrow$ RootSolverOrthoPolynom(f, Df, $x_{0}$, $n$, $\varepsilon$);
11:     end for
12:     return ${\mathbf{x}}^{\mathrm{gl}}$;
13:end function
3.8 Generating the Legendre polynomials and their derivatives
The general form of four typical different orthogonal polynomials (such as Legendre polynomials $P_{n}(x)$, Chebyshev polynomials $T_{n}(x)$, Laguerre polynomials $L_{n}(x)$ and Hermite polynomials $H_{n}(x)$), denoted by $Y_{n}(x)$, can be specified by the following iterative formulae
$\left\{\begin{split}Y_{n}(x)&=(A_{n}x+B_{n})Y_{n-1}(x)-C_{n}Y_{n-2}(x),\quad n\geq 2\\ Y_{0}(x)&=c_{0}\\ Y_{1}(x)&=c_{1}x+c_{2}\end{split}\right.$ (57)
where $c_{0},c_{1},c_{2}$ are constants, $A_{n},B_{n},C_{n}$ are functions depending on $n$. Taking the derivative on the two sides, we immediately have
$\left\{\begin{split}Y^{\prime}_{n}(x)&=A_{n}Y_{n-1}(x)+(A_{n}x+B_{n})Y^{\prime}_{n-1}(x)-C_{n}Y^{\prime}_{n-2}(x)\\ Y^{\prime}_{0}(x)&=0\\ Y^{\prime}_{1}(x)&=c_{1}\end{split}\right.$ (58)
where $n\geq 2$. Particularly, let
$\left\{\begin{split}{A}^{\mathrm{le}}_{n}&=\frac{2n-1}{n},{B}^{\mathrm{le}}_{n}=0,{C}^{\mathrm{le}}_{n}=\frac{n-1}{n}\\ {c}^{\mathrm{le}}_{0}&=1,{c}^{\mathrm{le}}_{1}=1,{c}^{\mathrm{le}}_{2}=0\end{split}\right.$ (59)
then for $n\geq 2$ we can deduce that
$\left\{\begin{split}P_{n}(x)&={A}^{\mathrm{le}}_{n}xP_{n-1}(x)-{C}^{\mathrm{le}}_{n}P_{n-2}(x)\\ P_{0}(x)&=1\\ P_{1}(x)&=x\end{split}\right.$ (60)
and
$\left\{\begin{split}P^{\prime}_{n}(x)&={A}^{\mathrm{le}}_{n}P_{n-1}(x)+{A}^{\mathrm{le}}_{n}xP^{\prime}_{n-1}(x)-{C}^{\mathrm{le}}_{n}P^{\prime}_{n-2}(x)\\ P^{\prime}_{0}(x)&=0\\ P^{\prime}_{1}(x)&=1\end{split}\right.$ (61)
This implies that we can compute the $n$-th order Legendre polynomial $P_{n}(x)$ and its derivative $P^{\prime}_{n}(x)$ iteratively with a strategy of top-down:
  • firstly, we design general procedures OrthoPolynom in Algorithm 3.8 and DerivOrthoPolynom in Algorithm 3.8 for generating $Y_{n}(x)$ and $Y^{\prime}_{n}(x)$ respectively with the iterative formulaes directly;
  • secondly, we design specific procedures — AnLegendre in Algorithm 3.8, BnLegendre in Algorithm 3.8 and CnLegendre in Algorithm 3.8 — to calculate $A(n),B(n)$ and $C(n)$ respectively and use pointer to functions to compute the Legendre polynomials $P_{n}(x)$ and $P^{\prime}_{n}(x)$.
The procedures involved in computing the Legendre polynomial $P_{n}(x)$ and its derivative $P^{\prime}_{n}(x)$ are illustrated in Figure 10 and Figure 11 respectively.
{forest}
[PolynomLegendre [OrthoPolynom [A[AnLegendre]] [B[BnLegendre]] [C[CnLegendre]] ] ]
[…]
{forest}
[DerivPolynomLegendre [DerivOrthoPolynom [A[AnLegendre]] [B[BnLegendre]] [C[CnLegendre]] ] ]
[…]
 Algorithm 27 Compute the orthogonal polynomials $Y_{n}(x)$ according to the iterative formulae. (OrthoPolynom)
1:The variable $x$, order $n\in\mathbb{N}$, three constants$\left\{c_{0},c_{1},c_{2}\right\}$, three pointers to function $\texttt{A},\texttt{B},\texttt{C}:\mathbb{N}\to\mathbb{R}$.
2:The value of $Y_{n}(x)$
[…]
4:     $Y_{0}\leftarrow c_{0}$; // $Y_{0}(x)=c_{0}$;
5:     $Y_{1}\leftarrow c_{1}x+c_{2}$; // $Y_{1}(x)=c_{1}x+c_{2}$;
6:     $Y_{n}\leftarrow 0$; // Initialize $Y_{n}(x)$ with 00;
7:     if $(n==0)$ then
8:         $Y_{n}\leftarrow Y_{0}$;
9:     end if
10:     if $(n==1)$ then
11:         $Y_{n}\leftarrow Y_{1}$;
12:     end if
13:     for $k\in\left\{2,\cdots,n\right\}$ do
14:         $Y_{n}\leftarrow(\texttt{A}(k)\cdot x+\texttt{B}(k))\cdot Y_{1}-\texttt{C}(k)\cdot Y_{0}$;
15:         $Y_{0}\leftarrow Y_{1}$; // Update $Y_{k-2}(x)$ with $Y_{k-1}(x)$;
16:         $Y_{1}\leftarrow Y_{n}$; // Update $Y_{k-1}(x)$ with $Y_{k}(x)$;
17:     end for
18:     return $Y_{n}$;
19:end function
 Algorithm 28 Compute the derivative $Y^{\prime}_{n}(x)$ of the orthogonal polynomial $Y_{n}(x)$ according to the iterative formulae (DerivOrthoPolynom).
1:The variable $x$, order $n\in\mathbb{N}$, constants $c_{0},c_{1}$ and $c_{2}$, three pointers to function A, B and C.
2:The value of the derivative function $Y^{\prime}_{n}(x)$
3:function DerivOrthoPolynom($x$,$n$,$c_{0}$,$c_{1}$,$c_{2}$,A,B,C)
4:     $Y_{0}\leftarrow c_{0}$; // $Y_{0}(x)=c_{0}$
5:     $Y_{1}\leftarrow c_{1}x+c_{2}$; // $Y_{1}(x)=c_{1}x+c_{2}$
6:     $Y^{\prime}_{0}\leftarrow 0$; // $Y_{0}(x)=c_{1}\Longrightarrow Y^{\prime}_{0}(x)=0$
7:     $Y^{\prime}_{1}\leftarrow c_{1}$; // $Y_{1}(x)=c_{1}x+c_{2}\Longrightarrow Y^{\prime}_{1}(x)=c_{1}$
8:     $Y_{n}\leftarrow 0$; // Initialize $Y_{n}(x)$ with 00
9:     $Y^{\prime}_{n}\leftarrow 0$; // Initialize $Y^{\prime}_{n}(x)$ with 00;
10:     if $(n==0)$ then
11:         $Y_{n}\leftarrow Y_{0}$;
12:         $Y^{\prime}_{n}\leftarrow Y^{\prime}_{0}$;
13:     end if
14:     if $(n==1)$ then
15:         $Y_{n}\leftarrow Y_{1}$;
16:         $Y^{\prime}_{n}\leftarrow Y^{\prime}_{0}$;
17:     end if
18:     for $k\in\left\{2,\cdots,n\right\}$ do
19:         $Y_{n}\leftarrow(\texttt{A}(k)\cdot x+\texttt{B}(k))\cdot Y_{1}-\texttt{C}(k)\cdot Y_{0}$;
20:         $Y^{\prime}_{n}\leftarrow\texttt{A}(k)\cdot Y_{1}+(\texttt{A}(k)\cdot x+\texttt{B}(k))\cdot Y^{\prime}_{1}-\texttt{C}(k)\cdot Y^{\prime}_{0}$;
21:         $Y_{0}\leftarrow Y_{1}$; // Update $Y_{k-2}(x)$ with $Y_{k-1}(x)$;
22:         $Y_{1}\leftarrow Y_{n}$; // Update $Y_{k-1}(x)$ with $Y_{k}(x)$;
23:         $Y^{\prime}_{0}\leftarrow Y^{\prime}_{1}$; // Update $Y^{\prime}_{k-2}(x)$ with $Y^{\prime}_{k-1}(x)$;
24:         $Y^{\prime}_{1}\leftarrow Y^{\prime}_{n}$;// Update $Y^{\prime}_{k-1}(x)$ with $Y^{\prime}_{k}(x)$;
25:     end for
26:     return $Y^{\prime}_{n}$;
27:end function
 Algorithm 29 Compute the coefficient $A(n)$ in the iterative formulae for Legendre polynomial (AnLegendre).
1:Order $n\in\mathbb{N}$
2:The value of ${A}^{\mathrm{le}}_{n}$
3:function AnLegendre($n$)
4:     $A\leftarrow 2.0-1.0/n$;
5:     return $A$;
6:end function
 Algorithm 30 Compute the coefficient $B(n)$ in the iterative formulae for Legendre polynomial (BnLegendre).
1:Order $n\in\mathbb{N}$
2:The value of ${B}^{\mathrm{le}}_{n}$
3:function BnLegendre($n$)
4:     return 00;
5:end function
 Algorithm 31 Compute the coefficient $C(n)$ in the iterative formulae for Legendre polynomial (CnLegendre).
1:Order $n\in\mathbb{N}$
2:The value of ${C}^{\mathrm{le}}_{n}$
3:function CnLegendre($n$)
4:     $C\leftarrow 1.0-1.0/n$;
5:     return $C$;
6:end function
 Algorithm 32 Compute the Legendre Polynomial $P_{n}(x)$ (PolynomLegendre).
1:The variable $x$, order $n\in\mathbb{N}$
2:The value of $P_{n}(x)$
3:function PolynomLegendre($x,n$)
4:     $c_{0}\leftarrow 1,c_{1}\leftarrow 1,c_{2}\leftarrow 0$;
[…]
8:     $P_{n}\leftarrow\textsc{OrthoPolynom}(x,n,c_{0},c_{1},c_{2},\texttt{A},\texttt{B},\texttt{C})$;
9:     return $P_{n}$;
10:end function
 Algorithm 33 Compute the derivative of Legendre polynomial (DerivPolynomLegendre).
1:The variable $x$, order $n\in\mathbb{N}$
2:The derivative of $P_{n}(x)$, i.e., the value of $P^{\prime}_{n}(x)$
3:function DerivPolynomLegendre($x,n$)
4:     $c_{0}\leftarrow 1$;
5:     $c_{1}\leftarrow 1$;
6:     $c_{2}\leftarrow 0$;
[…]
10:     $P^{\prime}_{n}\leftarrow\textsc{DerivOrthoPolynom}(x,n,c_{0},c_{1},c_{2},\texttt{A},\texttt{B},\texttt{C})$;
11:     return $P^{\prime}_{n}$;
12:end function
It should be remarked that designing is a mixture of multiple factors in a general viewpoint. Our discussion is limited in a narrow perspective which emphasizing algorithm design. For more deep thinking about design, please see Brooks’ famous book on design[11].
4 Verification, Validation and Testing
For the algorithms designed for computing the CEI-1, it is necessary for us to construct some simple but useful test cases for verification, validation and testing.
The verification of the four methods for solving CEI-1 is simple if each modules involved have been verified and validated. In Appendix D, we illustrate some typical test cases and omit the details for testing the modules for constructing the four numerical solvers for CEI-1 since it is simple and straight forward. After that, what we need do is just to compare the values of $K(k)$ obtained by Algorithm 3.1, Algorithm 3.2, Algorithm 3.3 and Algorithm 3.5.
Table 1: A comparison of numerical solvers for $K(k)$ and ${\mathcal{K}}_{\mathrm{cs}}(m)$ in MATLAB and Mathematica such that $m=k^{2}$
[…]
$0.00$ 1.5707963 $0.00$ 1.5707963
$0.10$ 1.5747456 $0.01$ 1.5747456
$0.20$ 1.5868678 $0.04$ 1.5868678
$0.30$ 1.6080486 $0.09$ 1.6080486
$0.40$ 1.6399999 $0.16$ 1.6399999
$0.50$ 1.6857504 $0.25$ 1.6857504
Table 1 shows the difference of numerical solvers. It is easy to find that
  • Our four numerical algorithms lead to the same output and the correctness can be verified by the Kseries since it is based on the expansion of formulae (14) which is right definitely and can be handled manually.
  • The available funcitons ellipticK in MATLAB and EllipticK in Mathematica have consistent numerical solutions, which implies that the algorithms adopted may be the same although the solutions are not acceptable.
It should be remarked that the built-in function ${\mathcal{K}}_{\mathrm{cs}}(m)$ in MATLAB and Mathematica are different from the $K(k)$ in our implementation although they are equivalent essentially.
Table 2: Validation of CEI-1 via Precision of Computation
Type of Solver Algorithm Procedure Default Value of Precision: Relative or Absolute Error Configuration of Parameters for Numerical Integration
IS-CEI-1 Algorithm 3.1 $\textsc{Kseries}(k,\varepsilon)$ Relative error $\varepsilon=10^{-9}$
AGM-CEI-1 Algorithm 3.2 $\textsc{Kagm}(k,\varepsilon)$ Absolute error $\varepsilon=10^{-9}$
GC-CEI-1 Algorithm 3.3 $\textsc{Kintgc}(k,n)$ Absolute error: $\left|{\mathcal{E}}^{\mathrm{gc}}_{n}({f}_{\mathrm{gc}},1,k)\right|<\varepsilon=10^{-9}$ $(k,n)\in\left\{(0.1,4),(0.2,5),(0.3,6),(0.4,7),(0.5,8)\right\}$
GL-CEI-1 Algorithm 3.5 $\textsc{Kintgl}(k,n)$ Absolute error: $\left|{\mathcal{E}}^{\mathrm{gl}}_{n}({f}_{\mathrm{gl}},0,\pi/2,k)\right|<\varepsilon=10^{-9}$ $(k,n)=(0.1,2)$
We remark that the validation of the four methods for computing $K(k)$ should be noted because there are different parameters which control the precision of computation. As shown in Table 2, it is not difficult to find the following features:
  • For the IS-CEI-1 solver, the meaning of relative precision $\varepsilon$ is clear, and the absolute error is also applicable.
  • For the AGM-CEI-1 solver, the meaning of absolute error is clear, and the relative error could be adopted easily.
  • For the GC-CEI-1 solver, $n$ is an implicit parameter for control the precision. Given the absolute error $\varepsilon$ and the parameter $k$, we should estimate the integer ${n}^{\mathrm{gc}}_{f}$ mentioned in Appendix A.1 such that
    $\left|{\mathcal{E}}^{\mathrm{gc}}_{n}(f,1,k)\right|<\varepsilon$ (62)
    for $n>{n}^{\mathrm{gc}}_{f}$.
  • For the GL-CEI-1 solver, $n$ is also an implicit parameter for control the precision. Given the absolute error $\varepsilon$ and the parameter $k$, we should estimate the integer ${n}^{\mathrm{gl}}_{f}$ mentioned in Appendix A.2 such that
    $\left|{\mathcal{E}}^{\mathrm{gl}}_{n}(f,0,\pi/2,k)\right|<\varepsilon$ (63)
    for $n>{n}^{\mathrm{gl}}_{f}$.
Why do we use the implicit parameter $n$ instead of the explicit parameter $\varepsilon$ to control the precision of computation in the GC-CEI-1 and GL-CEI-1 methods? The answer lies in the difficulty of estimating the upper bounds of $\left|{f}_{\mathrm{gc}}^{(2n)}(\xi,k)\right|$ and $\left|{f}_{\mathrm{gl}}^{(2n)}(\xi,k)\right|$ for $\xi\in(-1,1)$ and $k\in[0,\sqrt{2}/2]$. Consequently, it is difficult for us to design algorithms to estimate the ${n}^{\mathrm{gl}}_{f}$ and ${n}^{\mathrm{gc}}_{f}$ with the precision $\varepsilon$ directly. This feature gives us a good example to demonstrate the advantages and disadvantages of the methods and algorithms developed. It is valuable in the sense of education although it is a pity in the sense of engineering. For the readers who cares the controllability of the iteration numbers $n$ according to the precision $\varepsilon$ in the MP problem, we suggest the Newton-Costes method in numerical integration.
5 Conclusions
The numerical schemes for computing the CEI-1 is essential for its applications. The available numerical solutions obtained by MATLAB and Mathematica are different due to the different of definitions for $K(k)$ and ${\mathcal{K}}_{\mathrm{cs}}(m)$, which implies that we should pay attention to the verification, validation and testing stage in research, development and STEM education in order to avoid potential errors.
Theoretical analysis shows that the correctness of the numerical solution can be verified by the analytic formulae in the form of infinite series. As alternative choices, we proposed four feasible numerical schemes and designed the Kseries, Kagm, Kintgc and Kintgl solvers for CEI-1 to obtain consistent numerical results based on infinite series method, AGM method, Gauss-Chebyshev & Gauss-Legendre numerical integrations. The key algorithms developed have been verified, validated and tested.
Although our emphasis is put on the numerical algorithms for the CEI-1, the methods adopted can also be used for the complete elliptic integrals of the second and third kinds. Furthermore, the key modules designed for the four solvers are designed for more general problems so as to satisfy various applications as more as possible. Particularly, the iterative algorithms for generating orthogonal polynomials with iterative formulae $Y_{n}(x)=[A(n)x+B(n)]Y_{n-1}(x)-C(n)Y_{n-2}(x)$ and their derivatives can be taken to compute the Legendre, Chebyshev, Laguerre and Hermite polynomials which have wide applications in mathematics, physics and engineering.
Appendix A Error Estimation of Gaussian Numerical Integration
A.1 Gauss-Chebyshev Formulae and Approximation Error
The numerical integration was discussed by Zhang [9] and Atkinson [4].
Theorem 1.
For $f\in C^{2n}[-1,1]$, the integral ${\int^{1}_{-1}\frac{f(x)}{\sqrt{1-x^{2}}}\operatorname{d}x}$ can be computed by the Gauss-Chebyshev formulae
$\int^{1}_{-1}\frac{f(x)}{\sqrt{1-x^{2}}}\operatorname{d}x=\sum_{i=1}^{n}{w}^{\mathrm{gc}}_{i}f({x}^{\mathrm{gc}}_{i})+{\mathcal{E}}^{\mathrm{gc}}_{n}(f).$
where
${\mathcal{E}}^{\mathrm{gc}}_{n}(f)=\frac{2\pi}{2^{2n}(2n)!}f^{(2n)}(\eta),\quad\eta\in(-1,1)$ (64)
is the approximation error.
By changing the variable $x$ with $t=x/a$, we can obtain
$\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x=\int^{1}_{-1}\frac{f(at,k)}{\sqrt{1-t^{2}}}\operatorname{d}t.$ (65)
According to (65) and the $k$-th order derivative formulae
$\left[\frac{\operatorname{d}{}}{\operatorname{d}{x}}\right]^{k}f(\alpha x+\beta)=\alpha^{k}f^{(k)}(\alpha x+\beta),$ (66)
we can deduce the following corollary.
Corollary 2.
For $f(\cdot,k)\in C^{2n}[-1,1]$ and fixed parameter $k$, the integral ${\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x}$ can be computed with the extended Gauss-Chebyshev formulae
$\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x=\sum^{n}_{i=1}{w}^{\mathrm{gc}}_{i}f(a{x}^{\mathrm{gc}}_{i},k)+{\mathcal{E}}^{\mathrm{gl}}_{n}(f,a,k)$ (67)
where $\xi\in(-a,a)$ and
${\mathcal{E}}^{\mathrm{gc}}_{n}(f,a,k)=\frac{2\pi\cdot a^{2n}}{2^{2n}(2n)!}f^{(2n)}(\xi,k)$ (68)
is the approximation error.
Theoretically, if there is an upper bound $M^{f}_{n}(k)$ for $f^{(2n)}(\xi,k)$ such that $\left|f^{(2n)}(\xi,k)\right|\leq M_{n}(k)$ where $\xi\in(-a,a)$, then we can find an integer
${n}^{\mathrm{gc}}_{f}={n}^{\mathrm{gc}}_{f}(k,\epsilon,a)=\arg\min_{n\in\mathbb{N}}\left\{n:\frac{2\pi a^{2n}}{2^{2n}(2n)!}M^{f}_{n}(k)<\epsilon\right\}$ (69)
such that ${\mathcal{E}}^{\mathrm{gc}}_{n}(f,a,k)<\epsilon$ when $n>{n}^{\mathrm{gc}}_{f}$.
For illustration, if we take $f(x)=\cos x$, then $f^{(2n)}(x)=(-1)^{n}\cos x,f^{(2n)}(kx)=k^{2n}(-1)^{n}\cos kx$. Therefore, the approximation error of integral ${\int^{1}_{-1}\frac{\cos kx}{\sqrt{1-x^{2}}}\operatorname{d}x}=J_{0}(\left|k\right|),\quad k\in\mathbb{R}$ with the Gauss-Chebyshev formulae will be
${\mathcal{E}}^{\mathrm{gc}}_{n}(\cos,1,k)$ $=\frac{2\pi}{2^{2n}(2n)!}k^{2n}f^{(2n)}(\xi,k),\quad\xi\in(-1,1)$
$=\frac{2\pi}{2^{2n}(2n)!}k^{2n}(-1)^{n}\cos(k\xi)$
For the given $\epsilon>0$, the upper bound of the absolute approximate error $\left|{\mathcal{E}}^{\mathrm{gc}}_{n}(\cos,1,k)\right|<\epsilon$ means that $M^{\cos}_{n}(k)=k^{2n}$ and
$\frac{2\pi}{2^{2n}(2n)!}k^{2n}<\epsilon$
since $\left|(-1)^{n}\cos(k\xi)\right|\leq 1$. For a fixed $k$, we can set the parameter $n$ as
${n}^{\mathrm{gc}}_{\cos}=\arg\min_{i\in\mathbb{N}}\left\{i:\frac{2\pi\cdot k^{2i}}{2^{2i}(2i)!}<\epsilon\right\}$
Similarly, for $f(x)=\mathrm{e}^{x}$, we can deduce that
${n}^{\mathrm{gc}}_{\mathrm{e}}=\arg\min_{i\in\mathbb{N}}\left\{i:\frac{2\pi\cdot k^{2i}\mathrm{e}^{k}}{2^{2i}(2i)!}<\epsilon\right\}$
For $\epsilon=10^{-9}$, we have the $(k,n)$ pairs demonstrated in Table 3
Table 3: Comparison of ${n}^{\mathrm{gc}}_{\cos}$ and ${n}^{\mathrm{gc}}_{\mathrm{e}}$
$k$ 0.1 0.5 1.5 5 10 50 100 200 400
${n}^{\mathrm{gc}}_{\cos}$ 4 6 7 8 10 14 18 22 28
${n}^{\mathrm{gc}}_{\mathrm{e}}$ 5 6 7 9 12 26 42 69 117
In general, we can find the ${n}^{\mathrm{gc}}_{f}$ such that the absolute approximation error is less than $\epsilon$ for the fixed parameter $k$ if the upper bound of $\left|f^{(2n)}(\xi,k)\right|$ can be estimated properly.
A.2 Gauss-Legendre Formulae and Approximation Error
Theorem 3.
For $f\in C^{2n}[-1,1]$, the integral ${\int^{1}_{-1}f(x)\operatorname{d}x}$ can be computed with the Gauss-Legendre formulae
$\int^{1}_{-1}f(x)\operatorname{d}x=\sum^{n}_{i=1}{w}^{\mathrm{gl}}_{i}f({x}^{\mathrm{gl}}_{i})+{\mathcal{E}}^{\mathrm{gl}}_{n}(f)$
where
${\mathcal{E}}^{\mathrm{gl}}_{n}(f)=\frac{(n!)^{4}}{[(2n)!]^{3}}\cdot\frac{2^{2n+1}}{2n+1}\cdot f^{(2n)}(\eta),\quad\eta\in(-1,1)$ (70)
is the approximation error.
The standard interval $[-1,1]$ can be converted to a general finite interval $[a,b]$ by a linear transform $\psi:x\mapsto t=(b-a)x/2+(b+a)/2$, which implies that
$\int^{b}_{a}f(x)\operatorname{d}x=\frac{b-a}{2}\int^{1}_{-1}f\left(\frac{b+a+(b-a)x}{2}\right)\operatorname{d}x.$ (71)
According to (66), we can find that
$\left[\frac{\operatorname{d}{}}{\operatorname{d}{x}}\right]^{2n}f\left(\alpha x+\beta\right)=\alpha^{2n}f^{(2n)}\left(\alpha x+\beta\right).$ (72)
Let $\xi=[b+a+(b-a)\eta]/2$ where $\eta\in(-1,1)$, then $\xi\in(a,b)$. Substitute (71) and (72) into (70), we immediately have
$\begin{split}\int^{b}_{a}f(x)\operatorname{d}x&=\frac{b-a}{2}\sum^{n}_{i=1}{w}^{\mathrm{gl}}_{i}f\left(\frac{b+a+(b-a){x}^{\mathrm{gl}}_{i}}{2},k\right)\\ &+\left[\frac{b-a}{2}\right]^{2n+1}\cdot\frac{(n!)^{4}}{[(2n)!]^{3}}\cdot\frac{2^{2n+1}}{2n+1}f^{(2n)}(\xi)\end{split}$ (73)
from Theorem 3.
We now generalize the definite integral ${\int^{b}_{a}f(x)\operatorname{d}x}$ to the abstract function ${\int^{b}_{a}f(x,k)\operatorname{d}x}$. The deduction above implies the following corollary.
Corollary 4.
For $f(\cdot,k)\in C^{2n}[a,b]$ and fixed parameter $k$, the integral ${\int^{b}_{a}f(x,k)\operatorname{d}x}$ can be computed with the extended Gauss-Legendre formulae
$\begin{split}&\int^{b}_{a}f(x,k)\operatorname{d}x\\ &=\frac{b-a}{2}\sum^{n}_{i=1}{w}^{\mathrm{gl}}_{i}f\left(\frac{b+a+(b-a){x}^{\mathrm{gl}}_{i}}{2},k\right)\\ &\quad+{\mathcal{E}}^{\mathrm{gl}}_{n}(f,a,b,k)\end{split}$ (74)
where $\xi\in[a,b]$ and
${\mathcal{E}}^{\mathrm{gl}}_{n}(f,a,b,k)=\frac{(n!)^{4}}{[(2n)!]^{3}}\cdot\frac{(b-a)^{2n+1}}{2n+1}f^{(2n)}(\xi,k)$ (75)
is the approximation error.
Theoretically, if there is an upper bound $M^{f}_{n}(k)$ for $f^{(2n)}(\xi,k)$ such that $\left|f^{(2n)}(\xi,k)\right|\leq M^{f}_{n}(k)$ where $\xi\in(a,b)$, then we can find an integer
$\begin{split}{n}^{\mathrm{gl}}_{f}&={u}^{\mathrm{gl}}_{f}(k,\epsilon,a,b)\\ &=\arg\min_{n\in\mathbb{N}}\left\{n:\frac{(n!)^{4}}{[(2n)!]^{3}}\frac{(b-a)^{2n+1}}{2n+1}M^{f}_{n}(k)<\epsilon\right\}\end{split}$ (76)
such that ${\mathcal{E}}^{\mathrm{gl}}_{n}(f,a,b,k)<\epsilon$ when $n>{n}^{\mathrm{gl}}_{f}(k,\epsilon,a,b)$.
Appendix B Orthogonal Polynomials
Table 4 shows the four typical orthogonal polynomials and the corresponding expressions of $A(n),B(n),C(n),c_{0},c_{1},c_{2}$ for the Legendre polynomials $P_{n}(x)$, Chebyshev polynomials $T_{n}(x)$, Laguerre polynomials $L_{n}(x)$ and Hermite polynomials $H_{n}(x)$ of interest in wide applications.
Table 4: Parameters for the iterative formulae of orthogonal polynomials $Y_{n}(x)$ such that $Y_{0}(x)=c_{0},Y_{1}(x)=c_{1}x+c_{2}$ and $Y_{n}(x)=[A(n)x+B(n)]Y_{n-1}(x)-C(n)Y_{n-2}(x)$ for $n\geq 2$.
Type $\rho(x)$ $Y_{n}(x)$ Domain $A(n)$ $B(n)$ $C(n)$ $c_{0}$ $c_{1}$ $c_{2}$
Legendre $1$ $P_{n}(x)=\cfrac{1}{2^{n}n!}\cdot\cfrac{\operatorname{d}^{n}}{\operatorname{d}x^{n}}[(x^{2}-1)^{n}]$ $[-1,1]$ $2-\frac{1}{n}$ 00 $1-\frac{1}{n}$ $1$ $1$ 00
Chebyshev $\cfrac{1}{\sqrt{1-x^{2}}}$ $T_{n}(x)=\cos(n\arccos x)$ $[-1,1]$ $2$ 00 $1$ $1$ $2$ 00
Laguerre $\mathrm{e}^{-x}$ $L_{n}(x)=\cfrac{\mathrm{e}^{x}}{n!}\cdot\cfrac{\operatorname{d}^{n}}{\operatorname{d}x^{n}}\left[x^{n}\mathrm{e}^{-x}\right]$ $[0,+\infty)$ $-\frac{1}{n}$ $2-\frac{1}{n}$ $1-\frac{1}{n}$ $1$ $-1$ $1$
Hermite $\mathrm{e}^{-x^{2}}$ $H_{n}(x)=(-1)^{n}\mathrm{e}^{x^{2}}\cfrac{\operatorname{d}^{n}}{\operatorname{d}x^{n}}\left[\mathrm{e}^{-x^{2}}\right]$ (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) $2$ 00 $2(n-1)$ $1$ $2$ 00
Appendix C Nowton’s Method for Solving Root
Algorithm  C is used to calculate the root of the non-linear equation $f(x)=0$ with an initial value $x_{0}$ via Newton’s method (also named by Newton-Raphson method). Essentially, the Newton’s method for solving root is a special kind of the fixed-point algorithm in which the contractive mapping is specified by (49). Thus what we should do is to define the updating function ${\mathcal{A}}_{\mathrm{Newton}}$. Therefore, we first design a procedure NewtonUpdate for ${\mathcal{A}}_{\mathrm{Newton}}$, then we design the procedure RootSolverNewton by calling the procedure FixedPointSolver. For the purpose of solving the $n$ roots of Legendre polynomial $P_{n}(x)$ encapsulated by the procedure CalcRootsPolynomLegendre, we should set the argument (pointer fo function) f in NewtonUpdate with the procedure PolynomLegendre and call the procedure RootSolverNewton for $n$ times by a loop. Figure 9 shows the logic relation of the procedures involved for calculating the $n$ roots of $P_{n}(x)$.
 Algorithm 34 Calculate the root of the non-linear equation $f(x)=0$ with an initial value $x_{0}$ via Newton’s method (RootSolverNewton)
1:Pointers to function f for the function $f(x)$, initial value $x_{0}$, precision $\epsilon$.
2:the root of $f(x)=0$.
3:function RootSolverNewton(f, $x_{0}$, $\epsilon$)
[…]
7:     return root;
8:end function
Appendix D Test Cases for Verification and Validation
The optional test cases for verifying and validating the modules involved in constructing the key algorithms and corresponding procedures for computing the CEI-1 are shown in Table 5.
Table 5: Optional Test Cases for Computational Tasks
Computational Task Optional Test Cases Parameters Binding Results for Verification
${S(x)=\sum^{\infty}_{j=0}c_{j}x^{j}}$ ${\cos x=\sum^{\infty}_{j=0}(-1)^{j}\frac{x^{2j}}{(2j)!}}$ ${c_{2j}=\frac{(-1)^{j}}{(2j)!}}$ ${\cos\pi=-1,\dots}$
${\sin x=\sum^{\infty}_{j=0}(-1)^{j}\frac{x^{2j+1}}{(2j+1)!}}$ ${c_{2j+1}=\frac{(-1)^{j}}{(2j+1)!}}$ ${\sin\frac{\pi}{2}=1,\cdots}$
${\mathrm{e}^{x}=\sum^{\infty}_{j=0}\frac{x^{j}}{j!}}$ ${c_{j}=\frac{1}{j!}}$ ${\mathrm{e}^{1}=2.71828\cdots,\cdots}$
${\arctan x=\sum^{\infty}_{j=0}(-1)^{j}\frac{x^{2j+1}}{2j+1}}$ ${c_{2j+1}=\frac{(-1)^{j}}{2j+1}}$ ${\arctan 1=\frac{\pi}{4}}$
$P_{n}(x)$ $P_{0}(x)$ $n=0$ $P_{1}(x)=1$
$P_{1}(x)$ $n=1$ $P_{1}(x)=x$
$P_{2}(x)$ $n=2$ ${P_{2}(x)=(3x^{2}-1)/2}$
$P_{3}(x)$ $n=3$ ${P_{3}(x)=(5x^{3}-3x)/2}$
$P^{\prime}_{n}(x)$ $P^{\prime}_{0}(x)$ $n=0$ $P^{\prime}_{1}(x)=0$
$P^{\prime}_{1}(x)$ $n=1$ $P^{\prime}_{1}(x)=1$
$P^{\prime}_{2}(x)$ $n=2$ ${P^{\prime}_{2}(x)=3x}$
$P^{\prime}_{3}(x)$ $n=3$ ${P^{\prime}_{3}(x)=(15x^{2}-3)/2}$
$f(x,\alpha)=0$ ${f(x)=\prod^{n}_{i=1}(x-\gamma_{i})=0}$ $\gamma_{1}<\gamma_{2}<\cdots<\gamma_{n}$ $x\in\left\{\gamma_{1},\cdots,\gamma_{n}\right\}$
$P_{n}(x)=0$ $P_{1}(x)=0$ $n=1$ $x=0$
$P_{2}(x)=0$ $n=2$ $x\in\left\{-\frac{\sqrt{3}}{3},\frac{\sqrt{3}}{3}\right\}$
$P_{3}(x)=0$ $n=3$ $x\in\left\{-\sqrt{\frac{3}{5}},0,\sqrt{\frac{3}{5}}\right\}$
${{F}_{\mathrm{C}}(k)=\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x}$ ${\int^{a}_{-a}\frac{1}{\sqrt{1-x^{2}}}\operatorname{d}x}$ $f(x,k)=1$ ${2\arcsin a}$
${\int^{1}_{-1}\frac{\cos kx}{\sqrt{1-x^{2}}}\operatorname{d}x}$ $f(x)=\cos kx$ ${J_{0}(k)}$ (Bessel function)
${\int^{1}_{-1}\frac{\mathrm{e}^{kx}}{\sqrt{1-x^{2}}}\operatorname{d}x}$ $f(x)=\mathrm{e}^{kx}$ ${\pi\mathrm{e}^{kx}}$
${{F}_{\mathrm{L}}(k)=\int^{b}_{a}f(x,k)\operatorname{d}x}$ ${\int^{b}_{a}\cos(x+k)\operatorname{d}x}$ $f(x,k)=\cos(x+k)$ $\sin(k+b)-\sin(k+a)$
${\int^{b}_{a}\frac{(k+x)^{2}}{3}\operatorname{d}x}$ ${f(x,k)=\frac{(x+k)^{2}}{3}}$ $(k+b)^{3}-(k+a)^{3}$
${\int^{b}_{a}\mathrm{e}^{kx}\operatorname{d}x}$ ${f(x,k)=\mathrm{e}^{kx}}$ ${\frac{\mathrm{e}^{kb}-\mathrm{e}^{ka}}{k}}$
$\operatorname{AGM}(a,b)$ ${\operatorname{AGM}(2,3)}$ $a=2,b=3$ ${\frac{\pi}{2}/\int^{\frac{\pi}{2}}_{0}\frac{\operatorname{d}\theta}{\sqrt{2^{2}\cos^{2}\theta+3^{2}\sin^{2}\theta}}}$
${\frac{1}{\operatorname{AGM}(1+x,1-x)}}$ $x\in[0,1)$ ${\sum^{\infty}_{j=0}\left[\frac{(2j-1)!!}{(2j)!!}\right]^{2}x^{2j}}$
$\operatorname{AGM}(a,b)$ $a,b\in\mathbb{R}^{+}$ $\operatorname{AGM}(b,a)$
$\operatorname{AGM}(a,b)$ $a,b\in\mathbb{R}^{+}$ $\operatorname{AGM}\left(\cfrac{a+b}{2},\sqrt{ab}\right)$
Table 5 lists some optional test cases for the computational tasks in computing the CEI-1. We give some interpretations here:
  • For computing the infinite series $S(x)=\sum^{\infty}_{j=0}c_{m}x^{m}$, we can use some typical functions (say $\cos x,\sin x,\mathrm{e}^{x}$ and $\arctan x$) and their expansions of series to verify the numerical solutions obtained by Algorithm 3.1 by comparing their analytical solutions.
  • For generating the Legendre polynomial $P_{n}(x)$ and its derivative $P^{\prime}_{n}(x)$ , we can use the expressions of $P_{0}(x),P_{1}(x),P_{2}(x)$ and $P_{3}(x)$ to test whether Algorithm 3.8, Algorithm 3.8, Algorithm 3.8, Algorithm 3.8 and the corresponding small procedures involved work well.
  • For computing the root of nonlinear equation $f(x,\alpha)=0$ with Algorithm 3.2, Algorithm 3.2 and Algorithm 3.7, we can construct
    $f(x)=\prod^{n}_{i=1}(x-\gamma_{i})$
    to test the numerical solution to the analytical solution $x_{i}=\gamma_{i}$ for positive $n$ and $\gamma_{1}<\gamma_{2}<\cdots<\gamma_{n}$, say $\gamma_{i}=i$.
  • For calculating the $n$ roots of $P_{n}(x)$ with Newton’s method, we can use the analytical solution to $P_{1}(x)=x=0,P_{2}(x)=(3x^{2}-1)/2=0,P_{3}(x)=x(5x^{2}-3)/2=0$ to verify the numerical solution obtained by Algorithm 3.7, Algorithm 3.7 and Algorithm 3.7.
  • For computing the ${{F}_{\mathrm{C}}(k)=\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-k^{2}}}\operatorname{d}x}$ with Algorithm 3.3, we can put $f(x,k)\in\left\{1,\cos kx,\mathrm{e}^{kx}\right\}$ and compare the numerical solutions with their analytical counterparts
    $\left\{\begin{split}\int^{1}_{-1}\frac{1}{\sqrt{1-x^{2}}}\operatorname{d}x&=\pi\\ \int^{1}_{-1}\frac{\cos kx}{\sqrt{1-x^{2}}}\operatorname{d}x&=J_{0}(k)\\ \int^{1}_{-1}\frac{\mathrm{e}^{kx}}{\sqrt{1-x^{2}}}\operatorname{d}x&=\pi\mathrm{e}^{kx}\end{split}\right.$ (77)
    where $J_{0}(\cdot)$ is the 00-th order Bessel function of the first kind.
  • For solving the ${{F}_{\mathrm{L}}(k)=\int^{b}_{a}f(x,k)\operatorname{d}x}$ with Algorithm 3.5, we can set
    $f(x,k)\in\left\{\cos(x+k),\frac{1}{3}(x+k)^{2},\mathrm{e}^{kx}\right\}$
    to check the numerical solutions. Note that
    $\lim_{k\to 0}\int^{b}_{a}\mathrm{e}^{kx}\operatorname{d}x$ $=\lim_{k\to 0}\frac{\mathrm{e}^{kb}-\mathrm{e}^{ka}}{k}$
    $=b-a=\int^{b}_{a}\lim_{k\to 0}\mathrm{e}^{kx}\operatorname{d}x$
    by the Newton-Leibnitz formulae and the rule of L’Hospital in calculus. This result also can be verified with numerical integration.
  • The verification of Algorithm 3.2 for solving $\operatorname{AGM}(a,b)$ is challenging if compared with the other computational tasks since there is no simple analytical expression for it. However, as illustrated at the bottom of Table 5, we can check the value of $\operatorname{AGM}(a,b)$ through cross-verification by the symmetric-iterative property and the relation with infinite series and integral according to equations (20), (21), (22) and (23) since the integral and infinite series can be verified with available methods as mentioned above.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Code availability statement
The algorithms developed in this paper can be obtained at the GitHub site https://github.com/GrAbsRD/CEI-1/tree/main where a long table with filename Kvalues.txt for the values of $K(k)$ is provided and the reader can run the code to verify the algorithms and reuse the code according to his/her interests.
Data availability statement
Not applicable
Acknowledgment
This work was supported partly by the National Natural Science Foundation of China under grant number 62167003, and partly by the Hainan Provincial Education and Teaching Reform Project of Colleges and Universities under grant number Hnjg2021-37.

### v5

@@
 Abstract
-The complete elliptic integral of the first kind (CEI-1) plays in a significant role in mathematics, physics and engineering.
-There is no simple formula for its computation, thus numerical algorithms are essential for coping with practical problems involved.
+The complete elliptic integral of the first kind (CEI-1) plays a significant role in mathematics, physics and engineering.
+There is no simple formula for its computation, thus numerical algorithms are essential for coping with the practical problems involved.
 The commercial implementations for the numerical solutions, such as the functions ellipticK and EllipticK provided by MATLAB and Mathematica respectively, are based on ${\mathcal{K}}_{\mathrm{cs}}(m)$ instead of the usual form $K(k)$ such that ${\mathcal{K}}_{\mathrm{cs}}(k^{2})=K(k)$ and $m=k^{2}$.
@@
 The general algorithms for computing orthogonal polynomials developed are significant byproducts in the sense of STEM education and scientific computation.
-Keywords: Complete elliptic integral of the first kind (CEI-1); Orthogonal polynomials; Verification-Validation-Testing (VVT); Infinte series; Arithmetic-geometric mean (AGM); Gauss numerical integration
+Keywords: Complete elliptic integral of the first kind (CEI-1); STEM education; Algorithm design; Orthogonal polynomials; Verification-Validation-Testing (VVT); Infinte series; Arithmetic-geometric mean (AGM); Gauss numerical integration
 1 Introduction
@@
   • computing the CEI-1 with the arithmetic-geometric mean (AGM) method which converges rapidly and iterates simply;
-  • converting the CEI-1 into a finite sum with $12n-$ 1 order algebraic precision by taking $n$-th order orthogonal polynomials via Gauss numerical integration approach.
+  • converting the CEI-1 into a finite sum with the $12n-$ 1 order algebraic precision by taking the $n$-th order orthogonal polynomials via Gauss numerical integration approach.
 The Gauss numerical method can be expressed by[4]
@@
 $(-1)^{j}\binom{-r}{j}=\binom{r+j-1}{j},\quad r\in\mathbb{R},j\in\mathbb{N}$ (10)
-and integral formulae
+and integral formula
 $\int^{\pi/2}_{0}\sin^{2j}u\operatorname{d}u=\frac{(2j-1)!!}{(2j)!!}\cdot\frac{\pi}{2},\quad j\in\mathbb{N}$ (11)
 where
-$\left\{\begin{split}(2j-1)!!&=(2j-1)(2j-3)\cdots 3\cdot 1\\ (2j)!!&=(2j)(2j-2)\cdots 4\cdot 2\\ 0!!&=(-1)!!=1\end{split}\right.$ (12)
-are the double factorial. Therefore, we can find the infinite series for the CEI-1, viz.,
+$\left\{\begin{split}&(2j-1)!!=(2j-1)(2j-3)\cdots 3\cdot 1\\ &(2j)!!=(2j)(2j-2)\cdots 4\cdot 2\\ &0!!=(-1)!!=1\end{split}\right.$ (12)
+are the fundamental results for the double factorial. Therefore, we can find the infinite series for the CEI-1, viz.,
 $K(k)=\frac{\pi}{2}\cdot\sum^{\infty}_{j=0}\left[\frac{(2j-1)!!}{(2j)!!}\right]^{2}k^{2j}$ (13)
@@
 For the 2-dim sequence $\left\{(u_{i},v_{i})\right\}^{\infty}_{j=0}$ such that
-$\left\{\begin{split}u_{i+1}&=\frac{u_{i}+v_{i}}{2},\quad i\geq 0\\ v_{i+1}&=\sqrt{u_{i}\cdot v_{i}},\quad i\geq 0\\ u_{0}&=a\\ v_{0}&=b\end{split}\right.$ (18)
+$\left\{\begin{split}u_{i+1}&=\frac{u_{i}+v_{i}}{2},\quad i\geq 0\\ v_{i+1}&=\sqrt{u_{i}\cdot v_{i}},\quad i\geq 0\\ u_{0}&=a\geq 0\\ v_{0}&=b\geq 0\end{split}\right.$ (18)
 it is easy to show that: both the limit of arithmetic mean $u_{i}$ and the limit of geometric mean $v_{i}$ exist, and the limits for $u_{i}$ and $v_{i}$ are equal.
@@
 hold obviously.
-There are lots of formulae about the AGM and the CEI-1, we just list the following necessary ones according to our requirements:
+There are lots of formula about the AGM and the CEI-1, we just list the following necessary ones according to our requirements:
 $\frac{1}{\operatorname{AGM}(1+x,1-x)}$ $=\sum^{\infty}_{j=0}\left[\frac{(2j-1)!!}{(2j)!!}\right]^{2}x^{2j}$ (22)
@@
 by the iterative property (21). Hence we have
-$K(k)=\frac{\pi/2}{\operatorname{AGM}(1+k,1-k)}=\frac{\pi/2}{\operatorname{AGM}(1,\sqrt{1-k^{2}})}$ (27)
+$K(k)=\frac{\pi/2}{\operatorname{AGM}(1,\sqrt{1-k^{2}})}$ (27)
 This is the famous AGM method proposed by King in 1924.
@@
 2.3 Gauss-Chebyshev Method
-According to Appendix A.1, for the abstract function
+According to the Appendix A.1, for the abstract function
 ${F}_{\mathrm{C}}(k)=\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x=\int^{1}_{-1}\frac{f(at,k)}{\sqrt{1-t^{2}}}\operatorname{d}t,$ (28)
-of the Chebyshev type defined by the integral of ${\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}}$ with parameter $k$, we can computed it by the extended Gauss-Chebyshev formulae of numerical integration.
+of the Chebyshev type defined by the integral of ${\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}}$ with parameter $k$, we can computed it by the extended Gauss-Chebyshev formula of numerical integration.
 Let ${x}^{\mathrm{gc}}_{i}$ denote the $i$-th root of $T_{n}(x)=\cos(n\arccos x)$, i.e.
@@
 ${F}_{\mathrm{L}}(k)=\int^{b}_{a}f(x,k)\operatorname{d}x=\frac{b-a}{2}\int^{1}_{-1}f\left(\frac{b+a}{2}+\frac{b-a}{2}t\right)\operatorname{d}t$ (38)
-defined by the integral of $f(x,k)$ with parameter $k$ such that $f(\cdot,k)\in C^{2n}[a,b]$ can be computed by the extended Gauss-Legendre formulae of numerical integration.
+defined by the integral of $f(x,k)$ with parameter $k$ such that $f(\cdot,k)\in C^{2n}[a,b]$ can be computed by the extended Gauss-Legendre formula of numerical integration.
 Let ${x}^{\mathrm{gl}}_{i}$ denote the $i$-th root of the $n$-th order Legendre polynomial $P_{n}(x)$ and ${w}^{\mathrm{gl}}_{i}$ denote the $i$-th weight corresponding to ${x}^{\mathrm{gl}}_{i}$ such that
@@
 The expression above for CEI-1 means that the essential work left for calculating $K(k)$ is to find the root-weight pairs for the Legendre polynomial $P_{n}(x)$.
-Although there are several equivalent definitions for the Legendre polynomial, the best way for computing the Legendre polynomial $P_{n}(x)$ is the following iterative formulae
+Although there are several equivalent definitions for the Legendre polynomial, the best way for computing the Legendre polynomial $P_{n}(x)$ is the following iterative formula
 $\left\{\begin{split}P_{n}(x)&=\frac{2n-1}{n}xP_{n-1}(x)-\frac{n-1}{n}P_{n-2}(x),~{}n\geq 2\\ P_{0}(x)&=1;\\ P_{1}(x)&=x.\end{split}\right.$ (46)
-This formulae implies that we can compute $P_{n}(x)$ one by one with an iterative process.
+This formula implies that we can compute $P_{n}(x)$ one by one with an iterative process.
 The Newton’s method for solving root of a nonlinear equation $f(x,\alpha)=0$ with a constant parameter $\alpha$ is specified by the iterative process with the form
@@
   • calling the procedure for the infinite series and multiplying some constant if necessary.
-The procedures for calculating CEI-1 with infinite series method is shown in Figure 3, in which CoefFun is a procedure name (pointer to function or similar entity) used to calling the procedure of computing $c_{j}$.
+The procedures for calculating CEI-1 with infinite series method is shown in Figure 3, in which CoefFun is a procedure name (pointer to function or similar entity) used to calling the procedure for computing $c_{j}$.
 {forest}
@@
 may be large numbers if $j>10$ and the overflow problem may happen in some computer language if the type of data is integer.
-However, we can cope with the the value of fraction safely instead of dealing with its denominator and numerator separately.
+However, we can cope with the the value of the fraction safely instead of dealing with its denominator and numerator separately.
 Actually, we have
@@
 1:Pointer to function $\texttt{CoefFun}:\mathbb{N}\to\mathbb{R}$ for computing the coefficient $c_{j}$, variable $x$, and precision $\varepsilon$ with default value $\varepsilon=10^{-9}$.
-2:The sum of the infinite series
+2:The value of $\sum^{\infty}_{j=0}c_{j}x^{j}$
 3:function InfinteSeriesSolver($\texttt{CoefFun},x,\varepsilon$)
@@
 It should be noted that the function $\texttt{CoefFun}:\mathbb{N}\to\mathbb{R}$ is used as an argument for the procedure InfiniteSeries in Algorithm 3.1, which implies that the InfiniteSeries is a high order function and the pointer to function will be involved for the implementation with different computer programming languages.
-We remark that the pointer to function is a concept in the C programming language and its counterparts in Python and MATLAB are lambda expression and function handle respectively.
+We remark that the pointer to function is a concept in the C programming language and its counterparts in Lisp/Python/Java/C++ and MATLAB/Octave are lambda expression and function handle respectively.
 Now we can design the procedure for calculating the CEI-1 conveniently. Algorithm 3.1 describes the procedure Kseries for computing CEI-1 with the infinite series method.
@@
 Figure 4: Procedures for calculating CEI-1 with AGM method.
-Let $\mathbf{x}={[x_{1},x_{2}]}^{\textsf{T}}\in\mathbb{R}^{{2}\times{1}}$, then the AGM method can be converted to a 2-dim fixed-point problem with the iterative formulae
+Let $\mathbf{x}={[x_{1},x_{2}]}^{\textsf{T}}\in\mathbb{R}^{{2}\times{1}}$, then the AGM method can be converted to a 2-dim fixed-point problem with the iterative formula
 $\mathbf{x}_{i+1}={\mathcal{A}}_{\mathrm{agm}}(\mathbf{x}_{i}),\quad i=0,1,2,\cdots$ (54)
@@
 4:     $\texttt{sum}\leftarrow 0$ ;
-5:     for $i\in\left\{1,\cdots,n\right\}$ do
+5:     for $i\in\left\langle 1,\cdots,n\right\rangle$ do
 6:         $\texttt{sum}\leftarrow\texttt{sum}+(x_{i}-y_{i})^{2}$;
@@
 10:end function
-The iterative process of AGM implies a special kind of iteration of fixed-point, which can be computed in a generic way with the help of while-do or do-while loop.
+The iterative process of AGM implies a special kind of iteration for the fixed-point, which can be computed in a generic way with the help of while-do or do-while loop.
 The procedure FixedPointSolver, as shown in Algorithm 3.2, computes the $n$-dim fixed-point iteratively.
@@
  Algorithm 20 Calculate the vector ${\mathbf{v}}^{\mathrm{gl}}_{f}({\mathbf{x}}^{\mathrm{gl}},k,a,b)$ of function $f(x,k)$ at the x-nodes with the coordinates ${u}^{\mathrm{gl}}_{i}=\left[b+a+(b-a){x}^{\mathrm{gc}}_{i}\right]/2$ (CalcVectorVfGL).
-1:Pointer to function $\texttt{f}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, order $n$, $n$-dim root vector ${\mathbf{x}}^{\mathrm{gl}}$, variable $a$, variable $b$, and variable $k$.
+1:The pointer to function $\texttt{f}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, the order $n$, the $n$-dim root vector ${\mathbf{x}}^{\mathrm{gl}}$, the variable $a$, the variable $b$, and the variable $k$.
 2:Vector ${\mathbf{v}}^{\mathrm{gc}}_{f}({\mathbf{x}}^{\mathrm{gl}},k,a,b)$.
 3:function CalcVectorVfGL(f, $n$, ${\mathbf{x}}^{\mathrm{gc}}$, $k$, $a$, $b$)
-4:     for $i\in\left\{1,\cdots,n\right\}$ do  //${v}^{\mathrm{gl}}_{f,i}({\mathbf{x}}^{\mathrm{gl}},k,a,b)$;
+4:     for $i\in\left\langle 1,\cdots,n\right\rangle$ do  //${v}^{\mathrm{gl}}_{f,i}({\mathbf{x}}^{\mathrm{gl}},k,a,b)$;
 5:         ${u}^{\mathrm{gl}}\leftarrow(b+a+(b-a){x}^{\mathrm{gc}}_{i})/2$;
@@
  Algorithm 21 Calculate the CEI-1 with the Gauss-Legendre method (Kintgl)
-1:The variable $k\in[0,1)$, positive integer $n$ with default value $n=50$
+1:The variable $k\in[0,1)$, the positive integer $n$ with default value $n=50$
 2:The value of CEI-1
@@
 3:function CalcWeightsGL(${\mathbf{x}}^{\mathrm{gl}},n$)
-4:     for $i\in\left\{1,\cdots,n\right\}$ do
+4:     for $i\in\left\langle 1,\cdots,n\right\rangle$ do
 5:         $u\leftarrow\textsc{DerivPolynomLegendre}({x}^{\mathrm{gl}}_{i},n)$;
@@
 Figure 8 demonstrates the system of solving the root vector of $P_{n}(x)$.
-Usually in the books about special mathematical functions [8, 9, 10], the different kinds of orthogonal polynomials, such as the Legendre polynomials $P_{n}(x)$, Chebyshev polynomials $T_{n}(x)$ of the first or second kind, Laguerre polynomials $L_{n}(x)$ and Hermite polynomials $H_{n}(x)$, are treated separately, which hides the general iterative formulae (58).
+Usually in the books about special mathematical functions [8, 9, 10], the different kinds of orthogonal polynomials, such as the Legendre polynomials $P_{n}(x)$, Chebyshev polynomials $T_{n}(x)$ of the first or second kind, Laguerre polynomials $L_{n}(x)$ and Hermite polynomials $H_{n}(x)$, are treated separately, which hides the general iterative formula (58).
 Once the $i$-th root ${x}^{\mathrm{gl}}_{i}$ is obtained, the corresponding weight ${w}^{\mathrm{gl}}_{i}$ can be determined by (39).
 The steps for solving the $n$ roots of $P_{n}(x)$ lie in three issues:
-  • designing the update procedure NewtonUpdate in Algorithm 3.7 by using $P_{n}(x)$ with the iteration formulae in Newton’s method;
+  • designing the update procedure NewtonUpdate in Algorithm 3.7 by using $P_{n}(x)$ with the iteration formula in Newton’s method;
   • setting proper initial values ${x}^{\mathrm{gl}}_{i}[0]$ for the $n$ roots of $P_{n}(x)$ according to (50) and call the procedure CalcRootsLegendre in Algorithm 3.7;
@@
  Algorithm 26 Calculate the $n$-dim root vector ${\mathbf{x}}^{\mathrm{gl}}={[{x}^{\mathrm{gl}}_{1},\cdots,{x}^{\mathrm{gl}}_{n}]}^{\textsf{T}}$ of the $n$–th order Legendre polynomial $P_{n}(x)$ such that $P_{n}({x}^{\mathrm{gl}}_{i})=0$ for $1\leq i\leq n$ (CalcRootsLegendre).
-1:Positive integer $n$
+1:The positive integer $n$
 2:The root vector ${\mathbf{x}}^{\mathrm{gl}}$ of $P_{n}(x)$
@@
 […]
-8:     for $i\in\left\{1,2,\cdots,n\right\}$ do
+8:     for $i\in\left\langle 1,2,\cdots,n\right\rangle$ do
 9:         $x_{0}\leftarrow\cos\cfrac{\pi(i+0.75)}{n+0.25}$; //initial value;
@@
 3.8 Generating the Legendre polynomials and their derivatives
-The general form of four typical different orthogonal polynomials (such as Legendre polynomials $P_{n}(x)$, Chebyshev polynomials $T_{n}(x)$, Laguerre polynomials $L_{n}(x)$ and Hermite polynomials $H_{n}(x)$), denoted by $Y_{n}(x)$, can be specified by the following iterative formulae
+The general form of four typical different orthogonal polynomials (such as Legendre polynomials $P_{n}(x)$, Chebyshev polynomials $T_{n}(x)$, Laguerre polynomials $L_{n}(x)$ and Hermite polynomials $H_{n}(x)$), denoted by $Y_{n}(x)$, can be specified by the following iterative formula
 $\left\{\begin{split}Y_{n}(x)&=(A_{n}x+B_{n})Y_{n-1}(x)-C_{n}Y_{n-2}(x),\quad n\geq 2\\ Y_{0}(x)&=c_{0}\\ Y_{1}(x)&=c_{1}x+c_{2}\end{split}\right.$ (57)
@@
 This implies that we can compute the $n$-th order Legendre polynomial $P_{n}(x)$ and its derivative $P^{\prime}_{n}(x)$ iteratively with a strategy of top-down:
-  • firstly, we design general procedures OrthoPolynom in Algorithm 3.8 and DerivOrthoPolynom in Algorithm 3.8 for generating $Y_{n}(x)$ and $Y^{\prime}_{n}(x)$ respectively with the iterative formulaes directly;
+  • firstly, we design general procedures OrthoPolynom in Algorithm 3.8 and DerivOrthoPolynom in Algorithm 3.8 for generating $Y_{n}(x)$ and $Y^{\prime}_{n}(x)$ respectively with the iterative formulas directly;
   • secondly, we design specific procedures — AnLegendre in Algorithm 3.8, BnLegendre in Algorithm 3.8 and CnLegendre in Algorithm 3.8 — to calculate $A(n),B(n)$ and $C(n)$ respectively and use pointer to functions to compute the Legendre polynomials $P_{n}(x)$ and $P^{\prime}_{n}(x)$.
@@
 […]
- Algorithm 27 Compute the orthogonal polynomials $Y_{n}(x)$ according to the iterative formulae. (OrthoPolynom)
+ Algorithm 27 Compute the orthogonal polynomials $Y_{n}(x)$ according to the iterative formula. (OrthoPolynom)
 1:The variable $x$, order $n\in\mathbb{N}$, three constants$\left\{c_{0},c_{1},c_{2}\right\}$, three pointers to function $\texttt{A},\texttt{B},\texttt{C}:\mathbb{N}\to\mathbb{R}$.
@@
 12:     end if
-13:     for $k\in\left\{2,\cdots,n\right\}$ do
+13:     for $k\in\left\langle 2,\cdots,n\right\rangle$ do
 14:         $Y_{n}\leftarrow(\texttt{A}(k)\cdot x+\texttt{B}(k))\cdot Y_{1}-\texttt{C}(k)\cdot Y_{0}$;
@@
 19:end function
- Algorithm 28 Compute the derivative $Y^{\prime}_{n}(x)$ of the orthogonal polynomial $Y_{n}(x)$ according to the iterative formulae (DerivOrthoPolynom).
-1:The variable $x$, order $n\in\mathbb{N}$, constants $c_{0},c_{1}$ and $c_{2}$, three pointers to function A, B and C.
+ Algorithm 28 Compute the derivative $Y^{\prime}_{n}(x)$ of the orthogonal polynomial $Y_{n}(x)$ according to the iterative formula (DerivOrthoPolynom).
+1:The variable $x$, the order $n\in\mathbb{N}$, the constants $c_{0},c_{1}$ and $c_{2}$, the three pointers to function A, B and C.
 2:The value of the derivative function $Y^{\prime}_{n}(x)$
@@
 17:     end if
-18:     for $k\in\left\{2,\cdots,n\right\}$ do
+18:     for $k\in\left\langle 2,\cdots,n\right\rangle$ do
 19:         $Y_{n}\leftarrow(\texttt{A}(k)\cdot x+\texttt{B}(k))\cdot Y_{1}-\texttt{C}(k)\cdot Y_{0}$;
@@
 27:end function
- Algorithm 29 Compute the coefficient $A(n)$ in the iterative formulae for Legendre polynomial (AnLegendre).
+ Algorithm 29 Compute the coefficient $A(n)$ in the iterative formula for Legendre polynomial (AnLegendre).
 1:Order $n\in\mathbb{N}$
@@
 6:end function
- Algorithm 30 Compute the coefficient $B(n)$ in the iterative formulae for Legendre polynomial (BnLegendre).
+ Algorithm 30 Compute the coefficient $B(n)$ in the iterative formula for Legendre polynomial (BnLegendre).
 1:Order $n\in\mathbb{N}$
@@
 5:end function
- Algorithm 31 Compute the coefficient $C(n)$ in the iterative formulae for Legendre polynomial (CnLegendre).
+ Algorithm 31 Compute the coefficient $C(n)$ in the iterative formula for Legendre polynomial (CnLegendre).
 1:Order $n\in\mathbb{N}$
@@
 For the algorithms designed for computing the CEI-1, it is necessary for us to construct some simple but useful test cases for verification, validation and testing.
-The verification of the four methods for solving CEI-1 is simple if each modules involved have been verified and validated.
-In Appendix D, we illustrate some typical test cases and omit the details for testing the modules for constructing the four numerical solvers for CEI-1 since it is simple and straight forward.
+The verification of the four methods for solving CEI-1 is simple if all of the modules involved have been verified and validated.
+In the Appendix D, we illustrate some typical test cases and omit the details for testing the modules for constructing the four numerical solvers for CEI-1 since it is simple and straight forward.
 After that, what we need do is just to compare the values of $K(k)$ obtained by Algorithm 3.1, Algorithm 3.2, Algorithm 3.3 and Algorithm 3.5.
@@
 Table 1 shows the difference of numerical solvers. It is easy to find that
-  • Our four numerical algorithms lead to the same output and the correctness can be verified by the Kseries since it is based on the expansion of formulae (14) which is right definitely and can be handled manually.
-  • The available funcitons ellipticK in MATLAB and EllipticK in Mathematica have consistent numerical solutions, which implies that the algorithms adopted may be the same although the solutions are not acceptable.
+  • Our four numerical algorithms lead to the same output and the correctness can be verified by the Kseries since it is based on the expansion of formula (14) which is right definitely and can be handled manually.
+  • The available funcitons ellipticK in MATLAB and EllipticK in Mathematica have consistent numerical solutions, which implies that the algorithms adopted may be the same.
 It should be remarked that the built-in function ${\mathcal{K}}_{\mathrm{cs}}(m)$ in MATLAB and Mathematica are different from the $K(k)$ in our implementation although they are equivalent essentially.
@@
 It is valuable in the sense of education although it is a pity in the sense of engineering.
-For the readers who cares the controllability of the iteration numbers $n$ according to the precision $\varepsilon$ in the MP problem, we suggest the Newton-Costes method in numerical integration.
 5 Conclusions
@@
 The available numerical solutions obtained by MATLAB and Mathematica are different due to the different of definitions for $K(k)$ and ${\mathcal{K}}_{\mathrm{cs}}(m)$, which implies that we should pay attention to the verification, validation and testing stage in research, development and STEM education in order to avoid potential errors.
-Theoretical analysis shows that the correctness of the numerical solution can be verified by the analytic formulae in the form of infinite series.
+Theoretical analysis shows that the correctness of the numerical solution can be verified by the analytic formula in the form of infinite series.
 As alternative choices, we proposed four feasible numerical schemes and designed the Kseries, Kagm, Kintgc and Kintgl solvers for CEI-1 to obtain consistent numerical results based on infinite series method, AGM method, Gauss-Chebyshev & Gauss-Legendre numerical integrations.
@@
 Although our emphasis is put on the numerical algorithms for the CEI-1, the methods adopted can also be used for the complete elliptic integrals of the second and third kinds.
-Furthermore, the key modules designed for the four solvers are designed for more general problems so as to satisfy various applications as more as possible.
-Particularly, the iterative algorithms for generating orthogonal polynomials with iterative formulae $Y_{n}(x)=[A(n)x+B(n)]Y_{n-1}(x)-C(n)Y_{n-2}(x)$ and their derivatives can be taken to compute the Legendre, Chebyshev, Laguerre and Hermite polynomials which have wide applications in mathematics, physics and engineering.
+Furthermore, the key modules for the four solvers are designed for more general problems so as to satisfy various applications as much as possible.
+Particularly, the iterative algorithms for generating orthogonal polynomials with iterative formula $Y_{n}(x)=[A(n)x+B(n)]Y_{n-1}(x)-C(n)Y_{n-2}(x)$ and their derivatives can be taken to compute the Legendre, Chebyshev, Laguerre and Hermite polynomials which have wide applications in mathematics, physics and engineering.
 Appendix A Error Estimation of Gaussian Numerical Integration
-A.1 Gauss-Chebyshev Formulae and Approximation Error
+A.1 Gauss-Chebyshev Formula and Approximation Error
 The numerical integration was discussed by Zhang [9] and Atkinson [4].
 Theorem 1.
-For $f\in C^{2n}[-1,1]$, the integral ${\int^{1}_{-1}\frac{f(x)}{\sqrt{1-x^{2}}}\operatorname{d}x}$ can be computed by the Gauss-Chebyshev formulae
+For $f\in C^{2n}[-1,1]$, the integral ${\int^{1}_{-1}\frac{f(x)}{\sqrt{1-x^{2}}}\operatorname{d}x}$ can be computed by the Gauss-Chebyshev formula
 $\int^{1}_{-1}\frac{f(x)}{\sqrt{1-x^{2}}}\operatorname{d}x=\sum_{i=1}^{n}{w}^{\mathrm{gc}}_{i}f({x}^{\mathrm{gc}}_{i})+{\mathcal{E}}^{\mathrm{gc}}_{n}(f).$
@@
 $\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x=\int^{1}_{-1}\frac{f(at,k)}{\sqrt{1-t^{2}}}\operatorname{d}t.$ (65)
-According to (65) and the $k$-th order derivative formulae
+According to (65) and the $k$-th order derivative formula
 $\left[\frac{\operatorname{d}{}}{\operatorname{d}{x}}\right]^{k}f(\alpha x+\beta)=\alpha^{k}f^{(k)}(\alpha x+\beta),$ (66)
@@
 Corollary 2.
-For $f(\cdot,k)\in C^{2n}[-1,1]$ and fixed parameter $k$, the integral ${\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x}$ can be computed with the extended Gauss-Chebyshev formulae
+For $f(\cdot,k)\in C^{2n}[-1,1]$ and fixed parameter $k$, the integral ${\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x}$ can be computed with the extended Gauss-Chebyshev formula
 $\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x=\sum^{n}_{i=1}{w}^{\mathrm{gc}}_{i}f(a{x}^{\mathrm{gc}}_{i},k)+{\mathcal{E}}^{\mathrm{gl}}_{n}(f,a,k)$ (67)
@@
 For illustration, if we take $f(x)=\cos x$, then $f^{(2n)}(x)=(-1)^{n}\cos x,f^{(2n)}(kx)=k^{2n}(-1)^{n}\cos kx$.
-Therefore, the approximation error of integral ${\int^{1}_{-1}\frac{\cos kx}{\sqrt{1-x^{2}}}\operatorname{d}x}=J_{0}(\left|k\right|),\quad k\in\mathbb{R}$ with the Gauss-Chebyshev formulae will be
+Therefore, the approximation error of integral ${\int^{1}_{-1}\frac{\cos kx}{\sqrt{1-x^{2}}}\operatorname{d}x}=J_{0}(\left|k\right|),\quad k\in\mathbb{R}$ with the Gauss-Chebyshev formula will be
 ${\mathcal{E}}^{\mathrm{gc}}_{n}(\cos,1,k)$ $=\frac{2\pi}{2^{2n}(2n)!}k^{2n}f^{(2n)}(\xi,k),\quad\xi\in(-1,1)$
@@
 $\frac{2\pi}{2^{2n}(2n)!}k^{2n}<\epsilon$
-since $\left|(-1)^{n}\cos(k\xi)\right|\leq 1$. For a fixed $k$, we can set the parameter $n$ as
+since $\left|(-1)^{n}\cos(k\xi)\right|\leq 1$. For the fixed $k$, we can set the parameter $n$ as
 ${n}^{\mathrm{gc}}_{\cos}=\arg\min_{i\in\mathbb{N}}\left\{i:\frac{2\pi\cdot k^{2i}}{2^{2i}(2i)!}<\epsilon\right\}$
@@
 In general, we can find the ${n}^{\mathrm{gc}}_{f}$ such that the absolute approximation error is less than $\epsilon$ for the fixed parameter $k$ if the upper bound of $\left|f^{(2n)}(\xi,k)\right|$ can be estimated properly.
-A.2 Gauss-Legendre Formulae and Approximation Error
+A.2 Gauss-Legendre Formula and Approximation Error
 Theorem 3.
-For $f\in C^{2n}[-1,1]$, the integral ${\int^{1}_{-1}f(x)\operatorname{d}x}$ can be computed with the Gauss-Legendre formulae
+For $f\in C^{2n}[-1,1]$, the integral ${\int^{1}_{-1}f(x)\operatorname{d}x}$ can be computed with the Gauss-Legendre formula
 $\int^{1}_{-1}f(x)\operatorname{d}x=\sum^{n}_{i=1}{w}^{\mathrm{gl}}_{i}f({x}^{\mathrm{gl}}_{i})+{\mathcal{E}}^{\mathrm{gl}}_{n}(f)$
@@
 Corollary 4.
-For $f(\cdot,k)\in C^{2n}[a,b]$ and fixed parameter $k$, the integral ${\int^{b}_{a}f(x,k)\operatorname{d}x}$ can be computed with the extended Gauss-Legendre formulae
+For $f(\cdot,k)\in C^{2n}[a,b]$ and fixed parameter $k$, the integral ${\int^{b}_{a}f(x,k)\operatorname{d}x}$ can be computed with the extended Gauss-Legendre formula
 $\begin{split}&\int^{b}_{a}f(x,k)\operatorname{d}x\\ &=\frac{b-a}{2}\sum^{n}_{i=1}{w}^{\mathrm{gl}}_{i}f\left(\frac{b+a+(b-a){x}^{\mathrm{gl}}_{i}}{2},k\right)\\ &\quad+{\mathcal{E}}^{\mathrm{gl}}_{n}(f,a,b,k)\end{split}$ (74)
@@
 Table 4 shows the four typical orthogonal polynomials and the corresponding expressions of $A(n),B(n),C(n),c_{0},c_{1},c_{2}$ for the Legendre polynomials $P_{n}(x)$, Chebyshev polynomials $T_{n}(x)$, Laguerre polynomials $L_{n}(x)$ and Hermite polynomials $H_{n}(x)$ of interest in wide applications.
-Table 4: Parameters for the iterative formulae of orthogonal polynomials $Y_{n}(x)$ such that $Y_{0}(x)=c_{0},Y_{1}(x)=c_{1}x+c_{2}$ and $Y_{n}(x)=[A(n)x+B(n)]Y_{n-1}(x)-C(n)Y_{n-2}(x)$ for $n\geq 2$.
+Table 4: Parameters for the iterative formula of orthogonal polynomials $Y_{n}(x)$ such that $Y_{0}(x)=c_{0},Y_{1}(x)=c_{1}x+c_{2}$ and $Y_{n}(x)=[A(n)x+B(n)]Y_{n-1}(x)-C(n)Y_{n-2}(x)$ for $n\geq 2$.
 Type $\rho(x)$ $Y_{n}(x)$ Domain $A(n)$ $B(n)$ $C(n)$ $c_{0}$ $c_{1}$ $c_{2}$
@@
 ${\frac{1}{\operatorname{AGM}(1+x,1-x)}}$ $x\in[0,1)$ ${\sum^{\infty}_{j=0}\left[\frac{(2j-1)!!}{(2j)!!}\right]^{2}x^{2j}}$
-$\operatorname{AGM}(a,b)$ $a,b\in\mathbb{R}^{+}$ $\operatorname{AGM}(b,a)$
-$\operatorname{AGM}(a,b)$ $a,b\in\mathbb{R}^{+}$ $\operatorname{AGM}\left(\cfrac{a+b}{2},\sqrt{ab}\right)$
+$\operatorname{AGM}(a,b)$ $a,b\in\mathbb{R}^{+}$ $\operatorname{AGM}(b,a)=\operatorname{AGM}\left(\cfrac{a+b}{2},\sqrt{ab}\right)$
 Table 5 lists some optional test cases for the computational tasks in computing the CEI-1. We give some interpretations here:
@@
   • For generating the Legendre polynomial $P_{n}(x)$ and its derivative $P^{\prime}_{n}(x)$ , we can use the expressions of $P_{0}(x),P_{1}(x),P_{2}(x)$ and $P_{3}(x)$ to test whether Algorithm 3.8, Algorithm 3.8, Algorithm 3.8, Algorithm 3.8 and the corresponding small procedures involved work well.
-  • For computing the root of nonlinear equation $f(x,\alpha)=0$ with Algorithm 3.2, Algorithm 3.2 and Algorithm 3.7, we can construct
-    $f(x)=\prod^{n}_{i=1}(x-\gamma_{i})$
-    to test the numerical solution to the analytical solution $x_{i}=\gamma_{i}$ for positive $n$ and $\gamma_{1}<\gamma_{2}<\cdots<\gamma_{n}$, say $\gamma_{i}=i$.
+  • For computing the root of nonlinear equation $f(x,\alpha)=0$ with Algorithm 3.2, Algorithm 3.2 and Algorithm 3.7, we can construct $f(x)=\prod^{n}_{i=1}(x-\gamma_{i})$ to test the numerical solution to the analytical solution $x_{i}=\gamma_{i}$ for positive $n$ and $\gamma_{1}<\gamma_{2}<\cdots<\gamma_{n}$, say $\gamma_{i}=i$.
   • For calculating the $n$ roots of $P_{n}(x)$ with Newton’s method, we can use the analytical solution to $P_{1}(x)=x=0,P_{2}(x)=(3x^{2}-1)/2=0,P_{3}(x)=x(5x^{2}-3)/2=0$ to verify the numerical solution obtained by Algorithm 3.7, Algorithm 3.7 and Algorithm 3.7.
@@
     $=b-a=\int^{b}_{a}\lim_{k\to 0}\mathrm{e}^{kx}\operatorname{d}x$
-    by the Newton-Leibnitz formulae and the rule of L’Hospital in calculus. This result also can be verified with numerical integration.
+    by the Newton-Leibnitz formula and the rule of L’Hospital in calculus. This result also can be verified with numerical integration.
   • The verification of Algorithm 3.2 for solving $\operatorname{AGM}(a,b)$ is challenging if compared with the other computational tasks since there is no simple analytical expression for it.
@@
 The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
-Code availability statement
+Code & data availability statement
 The algorithms developed in this paper can be obtained at the GitHub site https://github.com/GrAbsRD/CEI-1/tree/main where a long table with filename Kvalues.txt for the values of $K(k)$ is provided and the reader can run the code to verify the algorithms and reuse the code according to his/her interests.
-Data availability statement
-Not applicable
 Acknowledgment